### v1

# Training-free Camera Control for Video Generation

Abstract
We propose a training-free and robust solution to offer camera movement control for off-the-shelf video diffusion models. Unlike previous work, our method does not require any supervised finetuning on camera-annotated datasets or self-supervised training via data augmentation. Instead, it can be plugged and played with most pretrained video diffusion models and generate camera controllable videos with a single image or text prompt as input. The inspiration of our work comes from the layout prior that intermediate latents hold towards generated results, thus rearranging noisy pixels in them will make output content reallocated as well. As camera move could also be seen as a kind of pixel rearrangement caused by perspective change, videos could be reorganized following specific camera motion if their noisy latents change accordingly. Established on this, we propose our method CamTrol, which enables robust camera control for video diffusion models. It is achieved by a two-stage process. First, we model image layout rearrangement through explicit camera movement in 3D point cloud space. Second, we generate videos with camera motion using layout prior of noisy latents formed by a series of rearranged images. Extensive experiments have demonstrated the robustness our method holds in controlling camera motion of generated videos. Furthermore, we show that our method can produce impressive results in generating 3D rotation videos with dynamic content. Project page at https://lifedecoder.github.io/CamTrol/.
1 Introduction
Figure 1: Training-free video camera control by CamTrol. CamTrol could handle basic camera motions, hybrid motions and complicated trajectories with precise camera coordinates. Besides, CamTrol produces impressive results on generating 3D rotation videos in various styles. Note that all human faces presented in this paper are synthesized using text-to-image models.
As a more appealing and content-richer modality, videos differ from images by including an extra temporal dimension. This temporal aspect provides increased versatility for depicting diverse and dynamic movements, which can be decomposed into object motion, background transitions and perspective changes. Recent years have witnessed the rapid development and splendid breakthrough of video generation with text prompt or images as input instructions li2023videogen ; hong2022cogvideo ; ho2022imagenvideo ; luo2023videofusion ; zeng2023PixelDance ; SVD ; sora ; ge2023preserve ; fei2023DysenVDM , and demonstrated the inestimable potential of diffusion models to synthesis realistic videos. While these video generation models have made progress in generating videos with highly dynamic objects and backgrounds zeng2023PixelDance ; SVD ; li2023videogen , most of them fail to provide camera control for the generated videos.
The difficulty of controlling camera trajectory in videos primarily arises from two aspects. The initial challenge lies in the inadequacy of annotated data. Most video annotations lack of descriptions, especially precise descriptions of video’s camera movements. As a result, video generation models trained on these data often fail to interpret text prompts related to camera motions and generate correct outputs. One solution to mitigate the data insufficiency problem is to mimic videos with camera movements through simple data augmentation yang2024direct . However, these methods could only handle simple camera motions like zoom or truck, and have trouble in dealing with more complicated ones. The second challenge lies in the effort of additional finetuning required for controlling camera movements. As camera trajectories could be sophisticated, they sometimes cannot be accurately elaborated using naïve text prompts alone. Common solutions wang2023motionctrl ; he2024cameractrl proposed to embed camera parameters into diffusion models through learnable encoders and perform extensive finetuning on large-scale datasets with detailed camera trajectories. However, such datasets like RealEstate10k realestate10k and MVImageNet yu2023mvimgnet are intensively limited in scale and diversity due to the difficulty associated with data collection, in this way, these finetuning methods demand substantial resources but exhibit limited generalizability to other types of data. Lack of annotations and heavy finetuning effort make camera control a challenging task in video generations.
In this work, we attempt to address these issues through a training-free solution to offer camera control for off-the-shelf video diffusion models. We begin by introducing two core observations underpin that video diffusion models can achieve camera movement control in a training-free manner. First, we find that base video models could produce results with rough camera moves by integrating specific camera-related text into input prompts, such as camera zooms in or camera pans right. This simple implementation, though not very accurate and always leads to static or wrong motions, shows the natural prior knowledge learnt by pretrained models about following different camera trajectories. The other observation is the effectiveness video models have exhibited in adapting to 3D generation tasks. Recent works sv3d ; im3d ; shi2023mvdream find that leveraging pretrained video models as initialization helps drastically improve the performance of multi-view generations, demonstrating their strong ability of handling perspective change. The two crucial observations reveals the hidden power of video models for camera motion control, thus, we seek to find a way to evoke this innate ability, as it already exists in the model itself.
We propose CamTrol, which offers camera control for off-the-shelf video diffusion models in a training-free but robust manner. CamTrol is inspired by the layout prior that noisy latents hold towards generation results: As pixels in noisy latents change their positions, corresponding rearrangement will also occur to the output and leads to layout modification. Considering camera moves could also be seen as a kind of layout rearrangement, this prior can serve as an efficacious hint providing video model with information of specific camera motions. Specifically, CamTrol consists of a two-stage procedure. In stage I, explicit camera movements are modeled in 3D point cloud representations and produce a series of rendered images indicating specific camera movements. In stage II, layout prior of noisy latents are utilized to guide video generations with camera movements. We conduct extensive experiments to validate the effectiveness of our proposed CamTrol. Both quantitative and qualitative results demonstrate the robustness of CamTrol as a useful tool of controlling camera motion for video diffusion models. Furthermore, we show that CamTrol  produces impressive results of dynamic 3D rotation videos in various styles.
2 Related Work
Camera Control for Video Generation
While methods aim for controlling video foundation models constantly emerge ma2024magic ; liu2023video ; feng2023ccedit , there are few works explore how to manipulate camera motion of generated videos. Earlier work hao2018controllable controls motion trajectory via warping image through densified sparse flow and pixel fusion, similar ideas also appear later in chen2023motion ; yin2023dragnuwa . Besides utilizing optical flow, two main techniques for implementing video camera control are via self-supervised augmentation or additional finetuning. yang2024direct disentangles object motion with camera movement and incorporates extra layers to embed camera motions, where model is trained in a self-supervised manner by augmenting input videos to stimulate simple camera movements. he2024cameractrl and wang2023motionctrl train an additional camera encoder and integrate the output into temporal attention layers of U-Net. guo2023animatediff learns new motion pattern via LoRA hu2021lora and finetuning with multiple reference videos.
Noise Prior of Latents in Diffusion Model
One of the most natural advantages of diffusion model comes from its pixel-wise noisy latents formed during denoising process. These latents hold strong causuality towards output and directly determine what the result looks like, meanwhile have robust error-resilience as they are perturbed by Gaussian noises across different scales. Numerous work have exploited the convenience of this noise prior to attain controllable generation, such as image-to-image translation meng2021sdedit , pixel-level manipulation nichol2022glide ; avrahami2022blended , image inpainting lugmayr2022repaint and semantic editing choi2021ilvr ; kim2022diffusionclip ; hou2024high . Recent research has shown that even sampled from Gaussian distribution, the initial noise of diffusion process still have significantly influence to the layout of generated contents mao2023guided . In other work, noise prior is used to guarantee temporal consistency among video frames luo2023videofusion , or to trade-off between fidelity and diversity of image editing kim2022diffusionclip .
Video Model for 3D Generation
Similar to how most video generation models using the groundwork laid by image foundation models blattmann2023alignVideo-LDM ; esser2023GEN1 ; singer2022make ; tuneavideo , training of 3D generation model also relies heavily on pretrained 2D video models sv3d ; im3d ; shi2023mvdream ; chen2024v3d ; han2024vfusion3d . These methods either finetunes with rendered videos directly SVD ; chen2024v3d ; im3d ; han2024vfusion3d , or adds camera embedding for each view as extra condition sv3d ; shi2023mvdream . Video foundation models have shown to be particularly beneficial in generating consistent multi-view rendering of 3D objects, demonstrates their inherent abundant prior knowledge for handling camera pose change.
3 Training-free Camera Control for Video Generation
CamTrol takes two stages to evoke the innate camera control ability hidden in base video models. In Sec. 3.1, we will describe how to model explicit camera motion in point cloud space. In Sec. 3.2, we will elaborate on the camera-controllable video generation with the guidance of noise layout prior.
Figure 2: Pipeline of CamTrol. In stage I, camera movements are modeled through explicit 3D point cloud. In stage II, layout prior of noisy latents are utilized to guide video generation.
3.1 Camera Motion Modeling
To evoke pretrained video diffusion model’s ability of dealing with camera perspective changes, hints of camera motion should be injected to diffusion model in a proper way. While simply concatenating camera trajectories with text prompt is incomprehensible for original model, previous works wang2023motionctrl ; he2024cameractrl introduce additional embedder to encode camera parameters and finetune with limited annotated data realestate10k ; yu2023mvimgnet , which are data-hungry yet lack of generalization ability. Other methods yang2024direct construct camera motions by self-supervised augmentations, but could only handle a few easy camera controls. Thus, we seek a more efficient and robust way to guide the model towards camera controllable.
Considering perspective change of video is originally caused by camera movements in 3D space, we resort to 3D representation for providing explicit motion hints to video diffusion models. Specifically, we choose point cloud as the intermediate representation, in which space we can expediently manipulate camera poses and positions for simulating diverse camera movements. Besides explicit camera modeling, introducing point cloud brings extra benefits: The first is its data-efficiency. By utilizing inpainting techniques, only one single input image is required for the whole point cloud reconstruction, this sidesteps the effort of large-scale finetuning. Second, consistency between multi-view renderings can be easily ensured, as the known points will remain unchanged once the reconstruction finished.
Point Cloud Initialization
We start by lifting the pixels in input image plane to 3D point cloud representations. In practice, the input image can be either user-defined or created by image generators like Stable Diffusion rombach2022high . Given an input image $\mathbf{I}_{0}\in\mathbb{R}^{3\times H\times W}$, we first estimate its depth map $\mathbf{D}_{0}$ using off-the-shelf monocular depth estimator ZeoDepth bhat2023zoedepth . By combining image and its depth map, point cloud $\mathcal{P}_{0}$ can be initialized as:
$\mathcal{P}_{0}=\phi([\mathbf{I}_{0},\mathbf{D}_{0}],\mathbf{K},\mathbf{P}_{0}),$ (1)
where $\phi$ denotes the mapping function from RGBD to 3D point cloud, $\mathbf{K}$ and $\mathbf{P}_{0}$ represent camera’s intrinsic and extrinsic matrices set by convention chung2023luciddreamer as they’re usually intractable.
Camera Trajectories
To get consistent images from multiple viewpoints, we set camera motion as a pre-defined trajectory of extrinsic matrix $\left\{\mathbf{P}_{1},...,\mathbf{P}_{N-1}\right\}$, each of which includes a rotation matrix and translation matrix representing camera’s pose and position. At each step $i$, we project the point cloud back to camera plane using function $\psi$ and get a rendered image with perspective change: $\mathbf{I}_{i}=\psi(\mathcal{P}_{i},\mathbf{K},\mathbf{P}_{i})$. By calculating extrinsic matrices of corresponding movement, we obtain a series of camera motions including zoom, tilt, pan, pedestal, truck, roll and rotate, enabling flexible camera movements. Detailed definitions of these movements are elaborated in Appendix A.1. By combining basic trajectories, hybrid camera movements can be attained and produce videos with cinematic charm. What’s more, benefit from explicit camera motion modeling, our method could support trajectories with precise coordinates, which means it can generate videos with any complicated camera motion (Fig. 1).
Multi-view Rendering
When perspective changes, there can be vacancies appear as some areas are unoccupied within the point cloud. To get more reasonable results, we employ image inpainting model rombach2022high to fill up the holes for new renderings, with a mask distinguishing the known points from nonexistent ones. After inpainted in 2D space, image is lifted again onto 3D space and gradually complete the whole point cloud representation. During this process, misalignment between adjacent views may occur since depth estimator only estimates relative depth, further leads to inconsistency of rendered images. We adopt depth coefficient optimization chung2023luciddreamer to avoid this situation, which can be formed as:
$d_{i}=\operatorname*{argmin}_{d}\left(\sum_{M}\left\|\phi([\tilde{\mathbf{I}}_{i},d\tilde{\mathbf{D}}_{i}],\mathbf{K},\mathbf{P}_{i})-\mathcal{P}_{i-1}\right\|\right),$ (2)
where $\tilde{\mathbf{I}}_{i}$ and $\tilde{\mathbf{D}}_{i}$ refer to the inpainted image and its depth map respectively, $d_{i}$ denotes depth coefficient to be optimized, and $M$ refers to the overlapping region between $\mathcal{P}_{i}$ and $\mathcal{P}_{i-1}$, as other areas are not shared for calculating $\ell_{1}$ loss.
Thus, we get a set of images refer to the input and indicate specific camera movement:
$\left\{\mathbf{I}_{0},...,\mathbf{I}_{N-1}\right\}=\left\{\psi(\mathcal{P}_{i},\mathbf{K},\mathbf{P}_{i})|i\in\left[0,N-1\right]\right\}.$ (3)
3.2 Layout Prior of Noise
With camera motion modeling, we obtain a sequence $\mathbf{V}_{0}=\left[\mathbf{I}_{0},...,\mathbf{I}_{N-1}\right]\in\mathbb{R}^{N\times 3\times H\times W}$ of rendered images adhering to a specific camera trajectory. Note that quality of rendered images are not perfect as single input image only leads to sparse point cloud reconstruction, besides, these renderings are static, thus they could not use directly as video frames. To form an ideal video, we need to find a way that satisfies the following three requirements: 1) camera motions should be maintained; 2) video should be encouraged with more dynamics; and 3) quality imperfection should be compensated.
Camera Motion Inversion
Recent work on diffusion models have demonstrated the strong controllability of its noisy latents meng2021sdedit ; mao2023guided , the causality and error-resilience they hold towards final output make them a convenient yet powerful tool for controllable generation of diffusion models. Particularly for initial noise, even sampled from Gaussian distribution, it still have significant influence on the layout of generated image mao2023guided , so that rearranging the noise pixels will make content in output relocate as well. For instance, if all pixels in initial noise shift to right by a certain distance, it is likely that generated output reflect a similar shift. This reminds us that the impact of camera movement on images could also be regarded as a kind of layout rearrangement, where pixels change their positions caused by viewpoint change. In a similar way, videos can be reorganized following camera motion if their noisy latents change accordingly.
Inspired by this, we first construct a series of noisy latents indicating corresponding camera movements. It can be intuitively done by employing diffusion model’s inversion process on rendered image sequence $\mathbf{V}_{0}$. Latent at timestep $t_{0}$ can be calculated as:
$\mathbf{V}_{t_{0}}=\sqrt{\bar{\alpha}_{t_{0}}}\mathbf{V}_{0}+\sqrt{1-\bar{\alpha}_{t_{0}}}\epsilon,\quad\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (4)
where $\bar{\alpha}_{t}$ are variances used in DDIM song2020DDIM scheduler. Because the rendered images $\mathbf{V}_{0}$ share common pixels in certain regions, their latents also have relevance to each other in a way indicating pixels’ move. Moreover, while being perturbed with random noise, blank spaces and flawed regions in $\mathbf{V}_{0}$ can be further filled with randomness, providing video model with more possibilities to generate and correct them.
Video Generation
After camera motion inversion, noisy latents presenting camera movements are then passed through the backward process of video diffusion model, utilizing their layout controllability to guide video generation. Leveraging prior knowledge of base video model, the generation process also bestows video with rational dynamic information. In this way, explicit camera movements are injected into video diffusion model in an appropriate and training-free fashion. Starting from noisy motion latents at timestep $t_{0}$, the generation step can be represented as:
$\hat{\mathbf{V}}_{t-1}=\sqrt{\alpha_{t-1}}\left(\frac{\mathbf{V}_{t}-\sqrt{1-\alpha_{t}}\epsilon^{(t)}_{\theta}(\mathbf{V}_{t})}{\sqrt{\alpha_{t}}}\right)+\sqrt{1-\alpha_{t-1}-\sigma^{2}_{t}}\epsilon^{(t)}_{\theta}(\mathbf{V}_{t})+\sigma_{t}\epsilon,\quad t\in\left[1,t_{0}\right].$ (5)
Here we use DDIM as sampling scheduler song2020DDIM and $\epsilon_{\theta}$ denotes the video model for noise prediction. $\sigma_{t}$ determines whether the denoising process is deterministic or probabilistic, we set $\sigma=1$ to encourage diversity of generation results.
Trade-off Between Fidelity and Diversity
Leveraging noise prior guidance in diffusion model could lead to trade-off problem between generation’s fidelity and diversity meng2021sdedit ; kim2022diffusionclip ; hou2024high , where results that hold more faithfulness towards guidance tend to decline in generation quality or diversity. In this task, similar circumstance also occurs as model is required to receive guidance from a series of imperfect renderings while generate a reasonable and dynamic video. The key factor to balance the trade-off problem lies in the choice of $t_{0}$. When larger $t_{0}$ applied, generation bears more resemblance to original guidance $\mathbf{V}_{0}$, yet lacks of rationality and dynamics to be an appealing video. Instead, smaller $t_{0}$ leads to well-generated video, but is less aligned with desired camera motion. In our experiments, we find larger $t_{0}$ works better for motions with moderate intensity, and for those with relatively drastic move, smaller $t_{0}$ shows preferable performance.
4 Experiments
4.1 Experimental Settings
Implementation Details
Our work is based on zeng2023PixelDance with 16 frames at resolution 256$\times$256, which is trained on public data. Note that our method could be incorporated with most video diffusion models to achieve camera control, thus we also include results performed on other backbones SVD ; luo2023videofusion . For fair comparison with other methods, we set all model’s DDIM steps as 25. Our method doesn’t require any additional training utilizing camera trajectories.
Evaluation Details
We compare CamTrol with state-of-the-art works: AnimateDiff guo2023animatediff , MotionCtrl wang2023motionctrl and CameraCtrl he2024cameractrl , all three methods finetune on specific data. For quantitative evaluation, FVD unterthiner2018FVD , FID heusel2017FID and IS saito2020IS are used to assess video generation quality, while CLIPSIM wu2021CLIPSIM quantifies the similarity between generated video and input prompt. The evaluations are performed on 8 basic camera trajectories, details could be found in Appendix A.1. For evaluation dataset, we follow similar setting in MotionCtrl wang2023motionctrl , but with more complex text prompt. Concretely, we apply UCF-101 soomro2012UCF as the reference of calculating FID and FVD, where we randomly sample 125 prompts per camera motion type and form 1000 samples in total for assessment. We also conduct a user study to compare different methods with human evaluation. Details can be found in Appendix A.5.
4.2 Comparisons with State-of-the-art Methods
Quantitative Evaluation
Quantitative evaluations are shown in Table  1. Compared to previous state-of-the-art work, CamTrol demonstrates superior performance across most quantitative metrics. As large-stride camera move sometimes causes deformation on objects, there appears little decrease in single frame’s quality, i.e. FID. While method like AnimateDiff guo2023animatediff exhibits better performance on this term, it takes as sacrifice the dynamics of generated video, which we will see in qualitative evaluations. For camera motion control, human evaluations from user study indicate that training-free CamTrol attains comparable performance even compared to other training-based methods, where the video diffusion models are heavily finetuned on specific camera motion data.
Table 1: Quantitative comparison. Experiments are conducted on 8 basic camera trajectories.
Method Metric User Study
[…]
AnimateDiff guo2023animatediff 3699.34 65.81 11.61 0.3073 1.12 0.82
MotionCtrl wang2023motionctrl 1768.25 78.26 14.60 0.3143 2.00 2.03
CameraCtrl he2024cameractrl 1383.68 76.18 14.04 0.3123 1.01 1.08
CamTrol 1134.71 69.70 15.87 0.3253 1.86 2.06
Figure 3: Qualitative comparison with finetune methods. CamTrol generates videos that adhere to desired camera movements, without compromising their dynamics and rationality.
Qualitative Analysis
Qualitative comparison of different methods is illustrated in Fig. 3. All methods generate videos align well with desired camera motion, however, the constraints imposed by camera trajectories are sometimes too strong, that output videos have to compromise their quality and rationality to satisfy these movements. Some examples that demonstrate this are results of Zoom Out and Roll Anticlockwise, which looks more like flat images being zoomed or rolled on 2D plane instead of videos recorded in 3D spaces. The reason might lies on the limitations of proprietary dataset used for finetuning realestate10k ; yu2023mvimgnet , where most scenes are static, making pretrained model forget about knowledge of dynamic contents. In contrast, CamTrol makes full use of the powerful dynamics in base model and generate plausible videos with camera perspective change utilizing layout prior guidance, handles these circumstances gracefully. More comparisons at Appendix A.3.
4.3 Ablation Study
Figure 4: Comparisons with base model. Controlling camera motion via prompt engineering doesn’t work at most times. Instead, CamTrol offers robust control towards video’s camera movement in a training-free manner.
Figure 5: Effectiveness of layout prior. Layout prior guidance compensates for the vacancies (Left) and static contents (Right) caused by point cloud rendering.
Comparison to Base Model
To demonstrate that changes of camera motion are attributed to our method rather than the innate capability of video base model, we conduct ablation study to assess its effectiveness. We add prompts describing relevant camera moves such as camera rotates around it or camera zooms out and let video model understand by itself. The results are shown in Fig. 4. We can observe that even provided with prompts indicating how camera should move, base model fail to produce correct results. Instead, CamTrol  is able to achieve corredponding motion control without any instructions from input prompts.
Effectiveness of Layout Prior
We employ ablation study to validate the effectiveness of layout prior guidance, and illustrate its necessity from two aspects: the completeness and dynamic of generated video. In Fig. 5, we showcase frames before and after noise prior guidance. With the position of camera changes, there appears regions unfilled in point cloud and causes blank spaces in rendered images (left part); Besides, due to the static nature of point cloud, rendered images remain stationary too (right part). Noise layout prior could compensate for these flaws, finally produce videos with inpainted vacancies and rationalized dynamics.
Effect of Timestep $t_{0}$
$t_{0}$ is a crucial factor that influences the trade-off between generated video’s diversity and its faithfulness to camera motions requirements. To investigate its effect on the output, we conduct experiments with various $t_{0}$ values, relevant results are shown in Fig. 7. As illustrated, videos generated with larger $t_{0}$ tend to conform better to camera motion requirements, but suffer from decrease in dynamics; On the contrary, smaller $t_{0}$ leads to more plausible generations but fails to meet camera’s requirements, as latents at these timesteps carry more randomness.
Generalization to Diverse Base Models
Our proposed CamTrol can be seamlessly plugged and played with most pretrained video diffusion models to achieve training-free camera control. We test its performance on different base video models, including SVD SVD and VideoFusion luo2023videofusion , the results are illustrated in Fig. 8. Our approach remains effective applied to other video base models, demonstrating its strong robustness and generalization ability. More results with other base models can be found in Appendix A.4.
4.4 Further Applications
Figure 6: Effect of $t_{0}$. Smaller $t_{0}$ encourages
dynamics while larger $t_{0}$ preserves camera
movements (Pedestal Down).
Figure 7: Camera control at different scales. CamTrol supports camera movements over various scales.
Figure 8: Applied onto SVD SVD and VideoFusion luo2023videofusion . CamTrol could be plugged and played with most video diffusion models to offer camera movement control.
3D Rotation Videos
One of the most advantages of our method is it can generate videos rotating around some objects and produce outputs similar to 3D generation models sv3d ; im3d . While these 3D models need large-scale training on 3D dataset and could only handle inputs in specific styles, our approach is able to deal with any type of images and achieve this in a completely zero-shot manner. Some results are shown in 1. We offer more results on multi-view synthesis in Appendix A.2.
Hybrid and Complex Camera Movements
By combining different basic camera trajectories, CamTrol can support hybrid camera movements and endow generated video with cinematic charm. Besides hybrid control, benefit from explicit camera motion modeling in 3D space, CamTrol could support trajectories with precise coordinates, which means it can generate videos with any complicated motion. Results about hybrid and complex camera motions are shown in Fig. 1 and Appendix A.2.
Camera Motion at Different Scales
CamTrol supports camera movements at controllable scales. We provide some results in Fig. 7. By specifying different magnitudes of camera’s extrinsic matrix within point cloud spaces, rendered images will exhibit varying degrees of motion, leading to videos with distinct scales of camera movements. This further demonstrate the powerful controllability of CamTrol, and provides new pathway for video’s customized camera control.
5 Conclusion
In this paper, we propose a training-free and robust method CamTrol to offer camera control for off-the-shelf video diffusion models. It consists of two-stage procedure including explicit camera motion modeling in 3D point cloud space and video generation utilizing layout prior of noisy latents. Compared to previours work, CamTrol  does not require any additional finetuning on camera-annotated datasets, or self-supervised training via data augmentation, instead, it could be plugged and played with most video diffusion models to generate camera controllable videos with single image or text prompt as input. Comprehensive experimental results demonstrate the effectiveness and robustness of CamTrol for controlling camera motion for videos. Besides, we show that CamTrol produces impressive results in generating videos at different scales, with 3D rotations and more complicated trajectories.
Appendix A Appendix
A.1 Definitions of Basic Camera Motions
We refer to the terminology in cinematography to describe different camera motions, definitions of each type are detailed in Table 2.
Table 2: Definitions of basic camera motions.
Camera Motion Directions Definition
Zoom In Camera moves towards or away from a subject.
Out
Tilt Up Rotating the camera vertically from a fixed position.
Down
Pan Left Rotating the camera horizontally from a fixed position.
Right
Pedestal Up Moving a camera vertically in its entirety.
Down
Truck Left Moving a camera horizontally in its entirety.
Right
Roll Clockwise Rotating a camera in its entirety in a horizontal manner.
Anticlockwise
Rotate Clockwise Moving a camera around a subject.
Anticlockwise
Hybrid Arbitrary Combination of other motions.
A.2 More Results on Camera Control
In this section, we showcase additional qualitative results of CamTrol on various camera movements. Fig. 9 shows video frames with camera move Rotate Anticlockwise and Rotate Clockwise. These outputs share sort of similarity with outputs of 3D generation models, as they all exhibit in a turning-table like way, which camera rotates around some objects. The difference here is that 3D model, as only trained on specific datasets, could only generate outputs in certain styles, e.g., single static object with no background. Instead, our model could handle arbitrary image as input, and generate a rotating-around video with proper dynamics. From this aspect, our method could be seen as a infinite source of attaining 3D data. And by utilizing our method with stronger backbones, video foundation models could truly become the largest source of 3D data as it should be.
Besides, we provide additional results on hybrid motions and basic motions, which is illustrated in Fig. 10, Fig. 11 and Fig. 13, respectively. We also present pre-loaded complex camera motions in Fig. 12.
A.3 More Comparisons to State-of-the-Art Methods
More comparison results with state-of-the-art methods are shown in Fig. 14 and Fig. 15. Compared to other works, our method matches well with input texts and camera requirements, meanwhile has wide-ranging diversities on generated contents.
A.4 More Results on other Base Video Models
Our methods could be plugged and played in arbitrary video diffusion models to achieve camera control. We test its performance on two other video generation backbones including SVD SVD and VideoFusion luo2023videofusion , relevant results are shown in Fig. 16.
A.5 Details of User Study
In user study, we collect 7 text prompts for each of the 8 basic camera trajectories, yielding 56 samples in total. Users are asked to sort randomly displayed four videos generated by four methods separately, and rate them from the highest score to the lowest score (belongs to {3,2,1}, where higher score indicates better performance). The assess metrics including two aspects: alignment of video move with input camera motion, and video quality (such as the deformation, rationality and diversity) of generated contents during camera moves. After rating, we calculate the average score attained by each methods.
The prompts applied in user study are ensured with balance between different categories. We involve prompts from various contents and styles including natural scenes, human, animations, objects, animals and etc.
Prompts applied in user study are listed as below:
  1. 1.
    Blueberries were squeezed by fingers.
  2. 2.
    The bear dolls were dancing in a circle, with lights flashing behind them, Christmas atmosphere.
  3. 3.
    The wind is blowing, and the flames are gradually extinguishing.
  4. 4.
    A young hero sets out on a journey to find their missing parent, encountering friends and foes along the way.
  5. 5.
    A toy robot discovers it can move on its own and sets off on an adventure around the house.
  6. 6.
    A determined ant leads its colony on a journey to find a new home.
  7. 7.
    The Great Wall of China winds its way across rugged mountains and through lush valleys, a testament to ancient engineering.
  8. 8.
    The bear doll in the basket is jumping out from the basket.
  9. 9.
    My armor pieces kept fluctuating and turned into scale-like armor.
  10. 10.
    A little boy turned a page of the book.
  11. 11.
    Jellyfish swim in the deep sea.
  12. 12.
    The big tree growing by the sea, its leaves slowly turn yellow and begin to wither and fall.
  13. 13.
    At the top of a mountain, a hiker is greeted by a breathtaking view of snow-capped peaks and rolling valleys below.
  14. 14.
    A group of squirrels gather around a picnic blanket, nibbling on sandwiches and stealing sips of lemonade.
  15. 15.
    The road is changing strangely and disappearing and twisting.
  16. 16.
    Clouds and fog linger on the tops of mountains.
  17. 17.
    Easter eggs showed different expressions, some happy, some confused, some sad, and some angry.
  18. 18.
    A skilled dancer twirls gracefully across the stage, captivating the audience with every movement.
  19. 19.
    A curious cat discovers a hidden treasure in the backyard.
  20. 20.
    At the beach, children build sandcastles while seagulls swoop overhead, and families relax under colorful umbrellas.
  21. 21.
    In a bustling marketplace, vendors call out their wares as shoppers weave through the crowds, browsing colorful stalls.
  22. 22.
    On the busy streets, people come and go.
  23. 23.
    Sunlight filters through dense foliage in a tranquil grove.
  24. 24.
    A toy monkey sits at a pretend computer, typing away furiously as it pretends to work on important documents.
  25. 25.
    The Eiffel Tower stands tall against a backdrop of Parisian rooftops, its iron lattice illuminated by the setting sun.
  26. 26.
    The Colosseum in Rome stands as a majestic reminder of ancient gladiator battles and Roman history.
  27. 27.
    A cozy reading nook by a window.
  28. 28.
    A regal statue of a historical figure stands tall in a city square, gazing out over the bustling streets.
  29. 29.
    Let them perform a dance.
  30. 30.
    The goose is playing trumpet. slow motion. 2D animation.
  31. 31.
    The machine is making Spider-Man’s hood.
  32. 32.
    The girl danced in the water, splashing water and her hair fluttering the wind.
  33. 33.
    The man looked up at the peak.
  34. 34.
    A woman reading and taking notes.
  35. 35.
    A group of penguins are walking towards you.
  36. 36.
    A shimmering portal appears in the midst of a moonlit forest, leading to realms unknown.
  37. 37.
    In a bustling port, cranes swing cargo containers onto waiting ships, orchestrating a dance of logistics and trade.
  38. 38.
    At dawn, the sky ignites in a riot of colors over the tranquil expanse of a mist-covered lake.
  39. 39.
    A touch of enchanted ink brings the pages of an ancient spellbook to life, casting spells with each turn.
  40. 40.
    A travel blogger documents their journey around the world, sharing tales of exotic destinations and unforgettable experiences.
  41. 41.
    A playful puppy learns to fetch a stick for the first time.
  42. 42.
    In a cozy living room, friends gather around a crackling fireplace, sipping hot cocoa and sharing stories on a chilly evening.
  43. 43.
    A man fell asleep amid the flowers, roses swaying in the wind.
  44. 44.
    A tutorial video guides viewers through the step-by-step process of creating intricate origami designs.
  45. 45.
    Underneath a starry sky, the Northern Lights dance in shimmering curtains of green and purple, painting the night with magic.
  46. 46.
    A cluttered desk in a home office, with papers strewn about and a computer humming softly.
  47. 47.
    A towering skyscraper gleams in the sunlight, its sleek glass facade reflecting the surrounding cityscape.
  48. 48.
    A modernist home with clean lines and geometric shapes, surrounded by minimalist landscaping.
  49. 49.
    A children’s playground is a kaleidoscope of colors, with slides, swings, and climbing structures painted in playful shades.
  50. 50.
    Thunder and rain, ancient oriental dragons hover in the air.
  51. 51.
    A dog barking at you.
  52. 52.
    Multiple people raised their arms and struggled upward in the lava of hell.
  53. 53.
    An astronaut aboard a space station encounters a mysterious anomaly while conducting experiments.
  54. 54.
    A guitarist strums their instrument with gentle fingers, the melody flowing naturally from their movements.
  55. 55.
    A snow-capped mountain peak towers above a tranquil alpine lake, mirrored perfectly in its glassy surface.
  56. 56.
    A rainbow arcs across the sky after a refreshing rain shower, its vibrant colors painting a stunning backdrop against the clouds.
Figure 9: Results with camera move Rotate Anticlockwise and Rotate Clockwise. Our method could generated 3D rotation-like videos with dynamics.
Figure 10: Results with camera move Hybrid I. Here we combine Zoom Out, Truck Left, Pedestal Up, Pan Right and Tilt Down together, form a cinematic like effect on generated videos.
Figure 11: Results with camera move Hybrid II. Here the hybrid motion is defined as Zoom In first, then Pedestal Up.
Figure 12: Results with complex camera motions.
Figure 13: Results with basic camera motions.
Figure 14: Comparison to state-of-the-art methods on Truck Left, Truck Right, Pedestal Up, Pedestal Down.
Figure 15: Comparison to state-of-the-art methods on Zoom In, Zoom Out, Roll Anticlockwise, Roll Clockwise.
Figure 16: Results with other video generation base models including SVD SVD and VideoFusion luo2023videofusion . Our method works well with these backbones.

### v2

@@
 Implementation Details
-Our work is based on zeng2023PixelDance with 16 frames at resolution 256$\times$256, which is trained on public data.
-Note that our method could be incorporated with most video diffusion models to achieve camera control, thus we also include results performed on other backbones SVD ; luo2023videofusion .
-For fair comparison with other methods, we set all model’s DDIM steps as 25.
+The major results presented in this paper are grounded on zeng2023PixelDance .
+To ensure a fair comparison with other state-of-the-art methods, we employ SVD as base model for quantitative evaluation.
+Inversion and generation steps among all methods are set to 25.
 Our method doesn’t require any additional training utilizing camera trajectories.
 Evaluation Details
-We compare CamTrol with state-of-the-art works: AnimateDiff guo2023animatediff , MotionCtrl wang2023motionctrl and CameraCtrl he2024cameractrl , all three methods finetune on specific data.
-For quantitative evaluation, FVD unterthiner2018FVD , FID heusel2017FID and IS saito2020IS are used to assess video generation quality, while CLIPSIM wu2021CLIPSIM quantifies the similarity between generated video and input prompt.
-The evaluations are performed on 8 basic camera trajectories, details could be found in Appendix A.1.
-For evaluation dataset, we follow similar setting in MotionCtrl wang2023motionctrl , but with more complex text prompt.
-Concretely, we apply UCF-101 soomro2012UCF as the reference of calculating FID and FVD, where we randomly sample 125 prompts per camera motion type and form 1000 samples in total for assessment.
-We also conduct a user study to compare different methods with human evaluation.
-Details can be found in Appendix A.5.
+We compare CamTrol with state-of-the-art works: AnimateDiff guo2023animatediff , MotionCtrl wang2023motionctrl and CameraCtrl he2024cameractrl , all three methods are finetune on SVD using specific data.
+In quantitative evaluation, FVD unterthiner2018FVD , FID heusel2017FID and IS saito2020IS are used to assess video generation quality, while CLIPSIM wu2021CLIPSIM quantifies the similarity between generated video and input prompt.
+With regard to the accuracy of camera motion, we adopt ParticleSFMzhao2022particlesfm and produce estimated camera trajectories from generated videos, with the use of Absolute Trajectory Error(ATE) measuring their differences compared to ground truth.
+Relative Pose Error(RPE) is calculated to assess between consecutive frames how well the relative motions match expected ones including their transition(RPE-T) and rotation part(RPE-R).
+Settings of evaluation dataset follow those established in MotionCtrl wang2023motionctrl , but include more complex trajectories.
+Specifically, we extract 51 distinct trajectories from RealEstate10krealestate10k , each paired with 10 prompts mentioned in MotionCtrl wang2023motionctrl , resulting in 510 samples in total for assessment.
+As AnimateDiff guo2023animatediff lacks the ability to handle complicated trajectories, it is only included in qualitative analysis part evaluating with 8 basic camera motions.
+Quantitative comparisons contain AnimateDiff guo2023animatediff can be found in Appendix A.3.
 4.2 Comparisons with State-of-the-art Methods
@@
 Quantitative evaluations are shown in Table  1.
-Compared to previous state-of-the-art work, CamTrol demonstrates superior performance across most quantitative metrics.
-As large-stride camera move sometimes causes deformation on objects, there appears little decrease in single frame’s quality, i.e.
-FID.
-While method like AnimateDiff guo2023animatediff exhibits better performance on this term, it takes as sacrifice the dynamics of generated video, which we will see in qualitative evaluations.
-For camera motion control, human evaluations from user study indicate that training-free CamTrol attains comparable performance even compared to other training-based methods, where the video diffusion models are heavily finetuned on specific camera motion data.
-Table 1: Quantitative comparison. Experiments are conducted on 8 basic camera trajectories.
-Method Metric User Study
-[…]
-AnimateDiff guo2023animatediff 3699.34 65.81 11.61 0.3073 1.12 0.82
-MotionCtrl wang2023motionctrl 1768.25 78.26 14.60 0.3143 2.00 2.03
-CameraCtrl he2024cameractrl 1383.68 76.18 14.04 0.3123 1.01 1.08
-CamTrol 1134.71 69.70 15.87 0.3253 1.86 2.06
-Figure 3: Qualitative comparison with finetune methods. CamTrol generates videos that adhere to desired camera movements, without compromising their dynamics and rationality.
+Building upon the same base model as MotionCtrl wang2023motionctrl and CameraCtrl he2024cameractrl , our method demonstrates superior performance across all quantitative metrics concerning both video quality and camera motion accuracy.
+Despite being training-free, CamTrol outperforms those that rely heavily on extensive finetuning and large-scale annotated data, attaining the lowest score in ATE, RPE-T and RPE-R.
+Benefit from explicit camera movements modeling and motion inversion, our method is capable to produce videos align best with those tricky trajectories while maintaining their visual qualities.
+Table 1: Quantitative comparisons.
+Method Video Quality Motion Accuracy
+FVD $\downarrow$ FID $\downarrow$ IS $\uparrow$ CLIP-SIM $\uparrow$ ATE $\downarrow$ RPE-T $\downarrow$ RPE-R $\downarrow$
+MotionCtrl wang2023motionctrl 3576.40 239.10 7.58 0.2933 4.006 1.086 0.106
+CameraCtrl he2024cameractrl 2922.99 243.98 8.07 0.2915 4.200 1.487 0.080
+SVD+CamTrol 2832.59 227.36 8.07 0.3100 3.917 0.947 0.017
+Figure 3: Qualitative comparisons with finetuned methods. CamTrol generates videos that adhere to desired camera movements, without compromising their dynamics and rationality.
 Qualitative Analysis
-Qualitative comparison of different methods is illustrated in Fig.
+Qualitative comparisons of different methods are illustrated in Fig.
 3.
@@
 4.3 Ablation Study
-Figure 4: Comparisons with base model.
+Figure 4: Comparison with base model.
 Controlling camera motion via prompt engineering doesn’t work at most times.
@@
 Comparison to Base Model
-To demonstrate that changes of camera motion are attributed to our method rather than the innate capability of video base model, we conduct ablation study to assess its effectiveness.
-We add prompts describing relevant camera moves such as camera rotates around it or camera zooms out and let video model understand by itself.
+To demonstrate that changes of camera motion are attributed to our method rather than the innate capability of video model, we conduct ablation study to assess its effectiveness.
+We add prompts describing certain camera moves(e.g.
+zooms out), letting video model understand by itself.
 The results are shown in Fig.
 4.
-We can observe that even provided with prompts indicating how camera should move, base model fail to produce correct results.
-Instead, CamTrol  is able to achieve corredponding motion control without any instructions from input prompts.
+It could be observed that even provided with prompts indicating how camera should move, base model fail to produce correct results.
+Instead, CamTrol  is able to implement designated motion control without any instructions from text prompts.
 Effectiveness of Layout Prior
-We employ ablation study to validate the effectiveness of layout prior guidance, and illustrate its necessity from two aspects: the completeness and dynamic of generated video.
+We employ ablation study to validate the effectiveness of layout prior guidance, illustrating its necessity from two aspects: the completeness of vacancies and dynamic of generated video.
 In Fig.
 5, we showcase frames before and after noise prior guidance.
-With the position of camera changes, there appears regions unfilled in point cloud and causes blank spaces in rendered images (left part); Besides, due to the static nature of point cloud, rendered images remain stationary too (right part).
+With camera pose changes, there appears regions unfilled in point cloud and causes blank spaces in rendered images(left part); Besides, due to the static nature of point cloud, rendered images remain stationary(right part).
 Noise layout prior could compensate for these flaws, finally produce videos with inpainted vacancies and rationalized dynamics.
@@
 Our proposed CamTrol can be seamlessly plugged and played with most pretrained video diffusion models to achieve training-free camera control.
-We test its performance on different base video models, including SVD SVD and VideoFusion luo2023videofusion , the results are illustrated in Fig.
+We present visual results showcasing its applications based on different video base models, including SVD SVD and VideoFusion luo2023videofusion , in Fig.
 8.
-Our approach remains effective applied to other video base models, demonstrating its strong robustness and generalization ability.
+Our approach remains effective applied to alternative video base models, demonstrating its strong robustness and generalization ability.
 More results with other base models can be found in Appendix A.5.
@@
 By combining different basic camera trajectories, CamTrol can support hybrid camera movements and endow generated video with cinematic charm.
-Besides hybrid control, benefit from explicit camera motion modeling in 3D space, CamTrol could support trajectories with precise coordinates, which means it can generate videos with any complicated motion.
-Results about hybrid and complex camera motions are shown in Fig.
+Besides this, explicit motion modeling also equips CamTrol with the abilities to support trajectories containing precise coordinates, which means it can generate videos presenting any complicated camera movement.
+Results about hybrid and complex motions are shown in Fig.
 1 and Appendix A.2.
@@
 By specifying different magnitudes of camera’s extrinsic matrix within point cloud spaces, rendered images will exhibit varying degrees of motion, leading to videos with distinct scales of camera movements.
-This further demonstrate the powerful controllability of CamTrol, and provides new pathway for video’s customized camera control.
+This further demonstrate the powerful controllability of CamTrol, and provides a new pathway for video’s customized camera control.
 5 Conclusion
@@
 Comprehensive experimental results demonstrate the effectiveness and robustness of CamTrol for controlling camera motion for videos.
-Besides, we show that CamTrol produces impressive results in generating videos at different scales, with 3D rotations and more complicated trajectories.
+Besides, we show that CamTrol produces impressive results in generating videos with 3D rotations, complicated trajectories, and different moving scales.
 Appendix A Appendix
@@
 Compared to other works, our method matches well with input texts and camera requirements, meanwhile has wide-ranging diversities on generated contents.
-A.4 More Results on other Base Video Models
-Our methods could be plugged and played in arbitrary video diffusion models to achieve camera control.
-We test its performance on two other video generation backbones including SVD SVD and VideoFusion luo2023videofusion , relevant results are shown in Fig.
+We also present quantitative comparisons that include AnimateDiff guo2023animatediff as part of the evaluation.
+As AnimateDiff guo2023animatediff lacks the ability to handle complicated trajectories, experiments are conducted focusing on 8 basic trajectories described in A.1.
+To expand the diversity of the models evaluated, we employ CamTrol on zeng2023PixelDance for assessment.
+With regard to basic trajectories, we apply UCF-101 soomro2012UCF as the reference of calculating FID and FVD, from where we randomly sample 125 prompts per camera motion type and form 1000 samples in total for assessment.
+Relevant results are exhibited in Table 3.
+Table 3: Quantitative comparisons on basic trajectories.
+Method Video Quality
+FVD $\downarrow$ FID $\downarrow$ IS $\uparrow$ CLIP-SIM $\uparrow$
+AnimateDiff guo2023animatediff 1582.93 73.79 13.39 0.3196
+MotionCtrl wang2023motionctrl 1768.25 78.26 14.60 0.3143
+CameraCtrl he2024cameractrl 1383.68 76.18 14.04 0.3123
+CamTrol+zeng2023PixelDance 1134.71 69.70 15.87 0.3253
+A.4 Ablation on Depth Optimization
+Ablation results relevant to the effectiveness of depth optimization are presented in Table 4.
+While it affects little on CLIPSIM, depth optimization exhibits benefits on other video quality assessments(FVD, FID, IS) and camera trajectory criteria(ATE, RPE).
+Experiments are conducted on complicated 10 trajectories extracted from RealEstate10k realestate10k , each with 10 prompts mentioned in MotionCtrl wang2023motionctrl .
+We assume that depth optimization offers better alignment between adjacent views rendered from point cloud, which benefits both perceptual quality and camera motion accuracy of generated videos.
+Table 4: Ablation on depth coefficient optimization.
+[…]
+w/o optimization 5033.44 219.01 5.67 0.2949 4.093 1.131 0.046
+w/ optimization 4977.88 218.36 5.72 0.2947 3.906 0.978 0.049
+A.5 More Results on other Base Video Models
+Our methods could be plugged and played with arbitrary video diffusion models to achieve camera control.
+We showcase its performance based on two alternative video generation backbones including SVD SVD and VideoFusion luo2023videofusion , relevant results are exhibited in Fig.
 16.
-A.5 Details of User Study
-In user study, we collect 7 text prompts for each of the 8 basic camera trajectories, yielding 56 samples in total.
-Users are asked to sort randomly displayed four videos generated by four methods separately, and rate them from the highest score to the lowest score (belongs to {3,2,1}, where higher score indicates better performance).
-The assess metrics including two aspects: alignment of video move with input camera motion, and video quality (such as the deformation, rationality and diversity) of generated contents during camera moves.
-After rating, we calculate the average score attained by each methods.
-The prompts applied in user study are ensured with balance between different categories.
-We involve prompts from various contents and styles including natural scenes, human, animations, objects, animals and etc.
-Prompts applied in user study are listed as below:
-  1. 1.
-    Blueberries were squeezed by fingers.
-  2. 2.
-    The bear dolls were dancing in a circle, with lights flashing behind them, Christmas atmosphere.
-  3. 3.
-    The wind is blowing, and the flames are gradually extinguishing.
-  4. 4.
-    A young hero sets out on a journey to find their missing parent, encountering friends and foes along the way.
-  5. 5.
-    A toy robot discovers it can move on its own and sets off on an adventure around the house.
-  6. 6.
-    A determined ant leads its colony on a journey to find a new home.
-  7. 7.
-    The Great Wall of China winds its way across rugged mountains and through lush valleys, a testament to ancient engineering.
-  8. 8.
-    The bear doll in the basket is jumping out from the basket.
-  9. 9.
-    My armor pieces kept fluctuating and turned into scale-like armor.
-  10. 10.
-    A little boy turned a page of the book.
-  11. 11.
-    Jellyfish swim in the deep sea.
-  12. 12.
-    The big tree growing by the sea, its leaves slowly turn yellow and begin to wither and fall.
-  13. 13.
-    At the top of a mountain, a hiker is greeted by a breathtaking view of snow-capped peaks and rolling valleys below.
-  14. 14.
-    A group of squirrels gather around a picnic blanket, nibbling on sandwiches and stealing sips of lemonade.
-  15. 15.
-    The road is changing strangely and disappearing and twisting.
-  16. 16.
-    Clouds and fog linger on the tops of mountains.
-  17. 17.
-    Easter eggs showed different expressions, some happy, some confused, some sad, and some angry.
-  18. 18.
-    A skilled dancer twirls gracefully across the stage, captivating the audience with every movement.
-  19. 19.
-    A curious cat discovers a hidden treasure in the backyard.
-  20. 20.
-    At the beach, children build sandcastles while seagulls swoop overhead, and families relax under colorful umbrellas.
-  21. 21.
-    In a bustling marketplace, vendors call out their wares as shoppers weave through the crowds, browsing colorful stalls.
-  22. 22.
-    On the busy streets, people come and go.
-  23. 23.
-    Sunlight filters through dense foliage in a tranquil grove.
-  24. 24.
-    A toy monkey sits at a pretend computer, typing away furiously as it pretends to work on important documents.
-  25. 25.
-    The Eiffel Tower stands tall against a backdrop of Parisian rooftops, its iron lattice illuminated by the setting sun.
-  26. 26.
-    The Colosseum in Rome stands as a majestic reminder of ancient gladiator battles and Roman history.
-  27. 27.
-    A cozy reading nook by a window.
-  28. 28.
-    A regal statue of a historical figure stands tall in a city square, gazing out over the bustling streets.
-  29. 29.
-    Let them perform a dance.
-  30. 30.
-    The goose is playing trumpet. slow motion. 2D animation.
-  31. 31.
-    The machine is making Spider-Man’s hood.
-  32. 32.
-    The girl danced in the water, splashing water and her hair fluttering the wind.
-  33. 33.
-    The man looked up at the peak.
-  34. 34.
-    A woman reading and taking notes.
-  35. 35.
-    A group of penguins are walking towards you.
-  36. 36.
-    A shimmering portal appears in the midst of a moonlit forest, leading to realms unknown.
-  37. 37.
-    In a bustling port, cranes swing cargo containers onto waiting ships, orchestrating a dance of logistics and trade.
-  38. 38.
-    At dawn, the sky ignites in a riot of colors over the tranquil expanse of a mist-covered lake.
-  39. 39.
-    A touch of enchanted ink brings the pages of an ancient spellbook to life, casting spells with each turn.
-  40. 40.
-    A travel blogger documents their journey around the world, sharing tales of exotic destinations and unforgettable experiences.
-  41. 41.
-    A playful puppy learns to fetch a stick for the first time.
-  42. 42.
-    In a cozy living room, friends gather around a crackling fireplace, sipping hot cocoa and sharing stories on a chilly evening.
-  43. 43.
-    A man fell asleep amid the flowers, roses swaying in the wind.
-  44. 44.
-    A tutorial video guides viewers through the step-by-step process of creating intricate origami designs.
-  45. 45.
-    Underneath a starry sky, the Northern Lights dance in shimmering curtains of green and purple, painting the night with magic.
-  46. 46.
-    A cluttered desk in a home office, with papers strewn about and a computer humming softly.
-  47. 47.
-    A towering skyscraper gleams in the sunlight, its sleek glass facade reflecting the surrounding cityscape.
-  48. 48.
-    A modernist home with clean lines and geometric shapes, surrounded by minimalist landscaping.
-  49. 49.
-    A children’s playground is a kaleidoscope of colors, with slides, swings, and climbing structures painted in playful shades.
-  50. 50.
-    Thunder and rain, ancient oriental dragons hover in the air.
-  51. 51.
-    A dog barking at you.
-  52. 52.
-    Multiple people raised their arms and struggled upward in the lava of hell.
-  53. 53.
-    An astronaut aboard a space station encounters a mysterious anomaly while conducting experiments.
-  54. 54.
-    A guitarist strums their instrument with gentle fingers, the melody flowing naturally from their movements.
-  55. 55.
-    A snow-capped mountain peak towers above a tranquil alpine lake, mirrored perfectly in its glassy surface.
-  56. 56.
-    A rainbow arcs across the sky after a refreshing rain shower, its vibrant colors painting a stunning backdrop against the clouds.
 Figure 9: Results with camera move Rotate Anticlockwise and Rotate Clockwise. Our method could generated 3D rotation-like videos with dynamics.

### v3

@@
 Second, we generate videos with camera motion using layout prior of noisy latents formed by a series of rearranged images.
-Extensive experiments have demonstrated the robustness our method holds in controlling camera motion of generated videos.
-Furthermore, we show that our method can produce impressive results in generating 3D rotation videos with dynamic content.
-Project page at https://lifedecoder.github.io/CamTrol/.
+Extensive experiments have demonstrated its superior performance in both video generation quality and camera motion alignment compared with other finetuned methods.
+Furthermore, we show the capacity of CamTrol generalizing to various base models, as well as its impressive applications in scalable motion control, dealing with complicated trajectories and unsupervised 3D video generation.
+Videos available at https://lifedecoder.github.io/CamTrol/.
 1 Introduction
-Figure 1: Training-free video camera control by CamTrol.
-CamTrol could handle basic camera motions, hybrid motions and complicated trajectories with precise camera coordinates.
-Besides, CamTrol produces impressive results on generating 3D rotation videos in various styles.
-Note that all human faces presented in this paper are synthesized using text-to-image models.
 As a more appealing and content-richer modality, videos differ from images by including an extra temporal dimension.
 This temporal aspect provides increased versatility for depicting diverse and dynamic movements, which can be decomposed into object motion, background transitions and perspective changes.
-Recent years have witnessed the rapid development and splendid breakthrough of video generation with text prompt or images as input instructions li2023videogen ; hong2022cogvideo ; ho2022imagenvideo ; luo2023videofusion ; zeng2023PixelDance ; SVD ; sora ; ge2023preserve ; fei2023DysenVDM , and demonstrated the inestimable potential of diffusion models to synthesis realistic videos.
-While these video generation models have made progress in generating videos with highly dynamic objects and backgrounds zeng2023PixelDance ; SVD ; li2023videogen , most of them fail to provide camera control for the generated videos.
+Recent years have witnessed the rapid development and splendid breakthrough of video generation with text prompt or images as input instructions (Li et al., 2023; Hong et al., 2022; Ho et al., 2022; Luo et al., 2023; Zeng et al., 2024; Blattmann et al., 2023a; Brooks et al., 2024; Ge et al., 2023; Fei et al., 2023), and demonstrated the inestimable potential of diffusion models to synthesis realistic videos.
+While these video generation models have made progress in generating videos with highly dynamic objects and backgrounds (Zeng et al., 2024; Blattmann et al., 2023a; Li et al., 2023), most of them fail to provide camera control for the generated videos.
 The difficulty of controlling camera trajectory in videos primarily arises from two aspects.
@@
 As a result, video generation models trained on these data often fail to interpret text prompts related to camera motions and generate correct outputs.
-One solution to mitigate the data insufficiency problem is to mimic videos with camera movements through simple data augmentation yang2024direct .
+One solution to mitigate the data insufficiency problem is to mimic videos with camera movements through simple data augmentation (Yang et al., 2024a).
 However, these methods could only handle simple camera motions like zoom or truck, and have trouble in dealing with more complicated ones.
@@
 As camera trajectories could be sophisticated, they sometimes cannot be accurately elaborated using naïve text prompts alone.
-Common solutions wang2023motionctrl ; he2024cameractrl proposed to embed camera parameters into diffusion models through learnable encoders and perform extensive finetuning on large-scale datasets with detailed camera trajectories.
-However, such datasets like RealEstate10k realestate10k and MVImageNet yu2023mvimgnet are intensively limited in scale and diversity due to the difficulty associated with data collection, in this way, these finetuning methods demand substantial resources but exhibit limited generalizability to other types of data.
+Common solutions (Wang et al., 2023; He et al., 2024) proposed to embed camera parameters into diffusion models through learnable encoders and perform extensive finetuning on large-scale datasets with detailed camera trajectories.
+However, such datasets like RealEstate10k (Zhou et al., 2018) and MVImageNet (Yu et al., 2023) are intensively limited in scale and diversity due to the difficulty associated with data collection, in this way, these finetuning methods demand substantial resources but exhibit limited generalizability to other types of data.
 Lack of annotations and heavy finetuning effort make camera control a challenging task in video generations.
@@
 The other observation is the effectiveness video models have exhibited in adapting to 3D generation tasks.
-Recent works sv3d ; im3d ; shi2023mvdream find that leveraging pretrained video models as initialization helps drastically improve the performance of multi-view generations, demonstrating their strong ability of handling perspective change.
+Recent works (Voleti et al., 2024; Melas-Kyriazi et al., 2024; Shi et al., 2023) find that leveraging pretrained video models as initialization helps drastically improve the performance of multi-view generations, demonstrating their strong ability of handling perspective change.
 The two crucial observations reveals the hidden power of video models for camera motion control, thus, we seek to find a way to evoke this innate ability, as it already exists in the model itself.
@@
 In stage II, layout prior of noisy latents are utilized to guide video generations with camera movements.
-We conduct extensive experiments to validate the effectiveness of our proposed CamTrol.
-Both quantitative and qualitative results demonstrate the robustness of CamTrol as a useful tool of controlling camera motion for video diffusion models.
-Furthermore, we show that CamTrol  produces impressive results of dynamic 3D rotation videos in various styles.
+Compared with previous works, CamTrol require no any additional finetuning utilizing camera annotated datasets, nor does it need self-supervised training based on data augmentation.
+Extensive experiments have demonstrated its superior performance in both video generation quality and camera motion alignment against other finetuned methods.
+Furthermore, we show the capacity of CamTrol generalizing to various base models, as well as its impressive applications in scalable motion control, dealing with complicated trajectories and unsupervised 3D video generation.
 2 Related Work
 Camera Control for Video Generation
-While methods aim for controlling video foundation models constantly emerge ma2024magic ; liu2023video ; feng2023ccedit , there are few works explore how to manipulate camera motion of generated videos.
-Earlier work hao2018controllable controls motion trajectory via warping image through densified sparse flow and pixel fusion, similar ideas also appear later in chen2023motion ; yin2023dragnuwa .
+While methods aim for controlling video foundation models constantly emerge (Ma et al., 2024; Liu et al., 2023; Feng et al., 2023), there are few works explore how to manipulate camera motion of generated videos.
+Earlier work (Hao et al., 2018) controls motion trajectory via warping image through densified sparse flow and pixel fusion, similar ideas also appear later in Chen et al.
+(2023) and Yin et al.
+(2023).
 Besides utilizing optical flow, two main techniques for implementing video camera control are via self-supervised augmentation or additional finetuning.
-yang2024direct disentangles object motion with camera movement and incorporates extra layers to embed camera motions, where model is trained in a self-supervised manner by augmenting input videos to stimulate simple camera movements.
-he2024cameractrl and wang2023motionctrl train an additional camera encoder and integrate the output into temporal attention layers of U-Net.
-guo2023animatediff learns new motion pattern via LoRA hu2021lora and finetuning with multiple reference videos.
+Yang et al.
+(2024a) disentangles object motion with camera movement and incorporates extra layers to embed camera motions, where model is trained in a self-supervised manner by augmenting input videos to stimulate simple camera movements.
+He et al.
+(2024) and Wang et al.
+(2023) train an additional camera encoder and integrate the output into temporal attention layers of U-Net.
+Guo et al.
+(2023) learns new motion pattern via LoRA (Hu et al., 2021) and finetuning with multiple reference videos.
 Noise Prior of Latents in Diffusion Model
@@
 These latents hold strong causuality towards output and directly determine what the result looks like, meanwhile have robust error-resilience as they are perturbed by Gaussian noises across different scales.
-Numerous work have exploited the convenience of this noise prior to attain controllable generation, such as image-to-image translation meng2021sdedit , pixel-level manipulation nichol2022glide ; avrahami2022blended , image inpainting lugmayr2022repaint and semantic editing choi2021ilvr ; kim2022diffusionclip ; hou2024high .
-Recent research has shown that even sampled from Gaussian distribution, the initial noise of diffusion process still have significantly influence to the layout of generated contents mao2023guided .
-In other work, noise prior is used to guarantee temporal consistency among video frames luo2023videofusion , or to trade-off between fidelity and diversity of image editing kim2022diffusionclip .
+Numerous work have exploited the convenience of this noise prior to attain controllable generation, such as image-to-image translation (Meng et al., 2021), pixel-level manipulation (Nichol et al., 2022), image inpainting (Lugmayr et al., 2022) and semantic editing (Choi et al., 2021; Hou et al., 2024).
+Recent research has shown that even sampled from Gaussian distribution, the initial noise of diffusion process still have significantly influence to the layout of generated contents (Mao et al., 2023).
+In other work, noise prior is used to guarantee temporal consistency among video frames (Luo et al., 2023), or to trade-off between fidelity and diversity of image editing (Kim et al., 2022).
 Video Model for 3D Generation
-Similar to how most video generation models using the groundwork laid by image foundation models blattmann2023alignVideo-LDM ; esser2023GEN1 ; singer2022make ; tuneavideo , training of 3D generation model also relies heavily on pretrained 2D video models sv3d ; im3d ; shi2023mvdream ; chen2024v3d ; han2024vfusion3d .
-These methods either finetunes with rendered videos directly SVD ; chen2024v3d ; im3d ; han2024vfusion3d , or adds camera embedding for each view as extra condition sv3d ; shi2023mvdream .
+Similar to how most video generation models using the groundwork laid by image foundation models (Blattmann et al., 2023b; Esser et al., 2023; Singer et al., 2022; Wu et al., 2023), training of 3D generation model also relies heavily on pretrained 2D video models (Voleti et al., 2024; Melas-Kyriazi et al., 2024; Shi et al., 2023; Chen et al., 2024; Han et al., 2024).
+These methods either finetunes with rendered videos directly (Blattmann et al., 2023a; Chen et al., 2024; Melas-Kyriazi et al., 2024; Han et al., 2024), or adds camera embedding for each view as extra condition (Voleti et al., 2024; Shi et al., 2023).
 Video foundation models have shown to be particularly beneficial in generating consistent multi-view rendering of 3D objects, demonstrates their inherent abundant prior knowledge for handling camera pose change.
 3 Training-free Camera Control for Video Generation
-CamTrol takes two stages to evoke the innate camera control ability hidden in base video models.
+Figure 1: Pipeline of CamTrol. In stage I, camera movements are modeled through explicit 3D point cloud . In stage II, layout prior of noisy latents are utilized to guide video generation.
+CamTrol takes two stages to evoke the innate camera control ability hidden in foundation models.
 In Sec.
-3.1, we will describe how to model explicit camera motion in point cloud space.
+3.1, we will describe how to model explicit camera movement for video generation.
 In Sec.
-3.2, we will elaborate on the camera-controllable video generation with the guidance of noise layout prior.
-Figure 2: Pipeline of CamTrol. In stage I, camera movements are modeled through explicit 3D point cloud . In stage II, layout prior of noisy latents are utilized to guide video generation.
+2, we will elaborate on video’s motion control with the guidance of noise layout prior.
 3.1 Camera Motion Modeling
 To evoke pretrained video diffusion model’s ability of dealing with camera perspective changes, hints of camera motion should be injected to diffusion model in a proper way.
-While simply concatenating camera trajectories with text prompt is incomprehensible for original model, previous works wang2023motionctrl ; he2024cameractrl introduce additional embedder to encode camera parameters and finetune with limited annotated data realestate10k ; yu2023mvimgnet , which are data-hungry yet lack of generalization ability.
-Other methods yang2024direct construct camera motions by self-supervised augmentations, but could only handle a few easy camera controls.
+While simply concatenating camera trajectories with text prompt is incomprehensible for original model, previous works (Wang et al., 2023; He et al., 2024) introduce additional embedder to encode camera parameters and finetune with limited annotated data (Zhou et al., 2018; Yu et al., 2023), which are data-hungry yet lack of generalization ability.
+Other method (Yang et al., 2024a) constructs camera motions by self-supervised augmentations, but could only handle a few easy camera controls.
 Thus, we seek a more efficient and robust way to guide the model towards camera controllable.
-Considering perspective change of video is originally caused by camera movements in 3D space, we resort to 3D representation for providing explicit motion hints to video diffusion models.
-Specifically, we choose point cloud as the intermediate representation, in which space we can expediently manipulate camera poses and positions for simulating diverse camera movements.
-Besides explicit camera modeling, introducing point cloud brings extra benefits: The first is its data-efficiency.
-By utilizing inpainting techniques, only one single input image is required for the whole point cloud reconstruction, this sidesteps the effort of large-scale finetuning.
-Second, consistency between multi-view renderings can be easily ensured, as the known points will remain unchanged once the reconstruction finished.
+Considering perspective change of video is originally caused by camera movements in 3D space, we resort to 3D representation for providing generation models with explicit motion hints.
+Specifically, we choose point cloud as the intermediate representation, in which space we can expediently manipulate camera poses for simulating diverse movements.
+One extra benefit that point cloud brings is its data-efficiency: By utilizing inpainting techniques, only one single input image is required for the whole point cloud reconstruction, this sidesteps the effort of large-scale finetuning.
 Point Cloud Initialization
 We start by lifting the pixels in input image plane to 3D point cloud representations.
-In practice, the input image can be either user-defined or created by image generators like Stable Diffusion rombach2022high .
-Given an input image $\mathbf{I}_{0}\in\mathbb{R}^{3\times H\times W}$, we first estimate its depth map $\mathbf{D}_{0}$ using off-the-shelf monocular depth estimator ZeoDepth bhat2023zoedepth .
+In practice, the input image can be either user-defined or created by image generators like Stable Diffusion (Rombach et al., 2022).
+Given an input image $\mathbf{I}_{0}\in\mathbb{R}^{3\times H\times W}$, we first estimate its depth map $\mathbf{D}_{0}$ using off-the-shelf monocular depth estimator ZeoDepth (Bhat et al., 2023).
 By combining image and its depth map, point cloud $\mathcal{P}_{0}$ can be initialized as:
 $\mathcal{P}_{0}=\phi([\mathbf{I}_{0},\mathbf{D}_{0}],\mathbf{K},\mathbf{P}_{0}),$ (1)
-where $\phi$ denotes the mapping function from RGBD to 3D point cloud, $\mathbf{K}$ and $\mathbf{P}_{0}$ represent camera’s intrinsic and extrinsic matrices set by convention chung2023luciddreamer as they’re usually intractable.
+where $\phi$ denotes the mapping function from RGBD to 3D point cloud, $\mathbf{K}$ and $\mathbf{P}_{0}$ represent camera’s intrinsic and extrinsic matrices set by convention (Chung et al., 2023) as they’re usually intractable.
 Camera Trajectories
-To get consistent images from multiple viewpoints, we set camera motion as a pre-defined trajectory of extrinsic matrix $\left\{\mathbf{P}_{1},...,\mathbf{P}_{N-1}\right\}$, each of which includes a rotation matrix and translation matrix representing camera’s pose and position.
+To get consistent images from multiple views, we set camera motion as a trajectory of extrinsic matrix $\left\{\mathbf{P}_{1},...,\mathbf{P}_{N-1}\right\}$, each including a rotation matrix and translation matrix representing camera’s pose and position.
 At each step $i$, we project the point cloud back to camera plane using function $\psi$ and get a rendered image with perspective change: $\mathbf{I}_{i}=\psi(\mathcal{P}_{i},\mathbf{K},\mathbf{P}_{i})$.
 By calculating extrinsic matrices of corresponding movement, we obtain a series of camera motions including zoom, tilt, pan, pedestal, truck, roll and rotate, enabling flexible camera movements.
-Detailed definitions of these movements are elaborated in Appendix A.1.
+Detailed definitions of these movements are elaborated in Appendix B.
 By combining basic trajectories, hybrid camera movements can be attained and produce videos with cinematic charm.
-What’s more, benefit from explicit camera motion modeling, our method could support trajectories with precise coordinates, which means it can generate videos with any complicated camera motion(Fig.
-1).
-Multi-view Rendering
+What’s more, benefit from explicit camera motion modeling, our method could support trajectories with precise extrinsics, which means it can generate videos with any complicated camera motion.
+Multi-view Consistency
 When perspective changes, there can be vacancies appear as some areas are unoccupied within the point cloud.
-To get more reasonable results, we employ image inpainting model rombach2022high to fill up the holes for new renderings, with a mask distinguishing the known points from nonexistent ones.
-After inpainted in 2D space, image is lifted again onto 3D space and gradually complete the whole point cloud representation.
-During this process, misalignment between adjacent views may occur since depth estimator only estimates relative depth, further leads to inconsistency of rendered images.
-We adopt depth coefficient optimization chung2023luciddreamer to avoid this situation, which can be formed as:
+To get more reasonable results, we employ image inpainting model (Rombach et al., 2022) to fill up the holes for new renderings, with a mask distinguishing the known points from nonexistent ones.
+This operation guarantees coherence between known views and novel views in 2D space.
+After inpainting, image is lifted again onto 3D space and gradually complete the whole point cloud representation.
+During this process, points between adjacent views may become misaligned since depth estimator only estimates relative depth, further leads to inconsistency in both 3D point cloud and rendered images.
+To avoid this situation, we adopt depth coefficient optimization (Chung et al., 2023) at each step of the camera movement, formed as:
 $d_{i}=\operatorname*{argmin}_{d}\left(\sum_{M}\left\|\phi([\tilde{\mathbf{I}}_{i},d\tilde{\mathbf{D}}_{i}],\mathbf{K},\mathbf{P}_{i})-\mathcal{P}_{i-1}\right\|\right),$ (2)
@@
 3.2 Layout Prior of Noise
-With camera motion modeling, we obtain a sequence $\mathbf{V}_{0}=\left[\mathbf{I}_{0},...,\mathbf{I}_{N-1}\right]\in\mathbb{R}^{N\times 3\times H\times W}$ of rendered images adhering to a specific camera trajectory.
-Note that quality of rendered images are not perfect as single input image only leads to sparse point cloud reconstruction, besides, these renderings are static, thus they could not use directly as video frames.
-To form an ideal video, we need to find a way that satisfies the following three requirements: 1) camera motions should be maintained; 2) video should be encouraged with more dynamics; and 3) quality imperfection should be compensated.
+Figure 2: Transition of samples between two distinct distributions.
+As the layout-arranged images being inverted by adding random noise, the distribution of their noisy latents will gradually converge to their video counterpart (green area), finally forming nearly Gaussian latents with specific layout change.
+This information is then inherited during video generation process.
+Steps of camera motion inversion determine the trade-off between video diversity and motion fidelity.
+With camera motion modeling, we obtain a sequence $\mathbf{V}_{0}=\left\{\mathbf{I}_{0},...,\mathbf{I}_{N-1}\right\}\in\mathbb{R}^{N\times 3\times H\times W}$ of rendered images adhering to a specific camera trajectory.
+Note that quality of rendered images are not perfect since single input image only leads to sparse point cloud reconstruction, besides, these renderings are static, thus they could not use directly as video frames.
+To form an ideal video, we need to find a way that satisfies three requirements: 1) camera motions should be maintained; 2) video should be encouraged with more dynamics; and 3) quality imperfection should be compensated.
 Camera Motion Inversion
-Recent work on diffusion models have demonstrated the strong controllability of its noisy latents meng2021sdedit ; mao2023guided , the causality and error-resilience they hold towards final output make them a convenient yet powerful tool for controllable generation of diffusion models.
-Particularly for initial noise, even sampled from Gaussian distribution, it still have significant influence on the layout of generated image mao2023guided , so that rearranging the noise pixels will make content in output relocate as well.
+Recent work on diffusion models have demonstrated the strong controllability of its noisy latents (Meng et al., 2021; Mao et al., 2023), the causality and error-resilience they hold towards final output make them a convenient yet powerful tool for controllable generation of diffusion models.
+Particularly for initial noise, even sampled from Gaussian distribution, it still have significant influence on the layout of generated image (Mao et al., 2023), so that rearranging the noise pixels will make content in output relocate as well.
 For instance, if all pixels in initial noise shift to right by a certain distance, it is likely that generated output reflect a similar shift.
@@
 In a similar way, videos can be reorganized following camera motion if their noisy latents change accordingly.
-Inspired by this, we first construct a series of noisy latents indicating corresponding camera movements.
-It can be intuitively done by employing diffusion model’s inversion process on rendered image sequence $\mathbf{V}_{0}$.
-Latent at timestep $t_{0}$ can be calculated as:
+Inspired by this, we first construct a series of noisy latents indicating specific camera movements.
+It can be intuitively done by employing diffusion’s inversion process on rendered image sequence $\mathbf{V}_{0}$.
+Latent at timestep $t_{0}$ can be calculated as follows, where $\bar{\alpha}_{t}$ are variances of the scheduler:
 $\mathbf{V}_{t_{0}}=\sqrt{\bar{\alpha}_{t_{0}}}\mathbf{V}_{0}+\sqrt{1-\bar{\alpha}_{t_{0}}}\epsilon,\quad\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (4)
-where $\bar{\alpha}_{t}$ are variances used in DDIM song2020DDIM scheduler.
 Because the rendered images $\mathbf{V}_{0}$ share common pixels in certain regions, their latents also have relevance to each other in a way indicating pixels’ move.
@@
 $\hat{\mathbf{V}}_{t-1}=\sqrt{\alpha_{t-1}}\left(\frac{\mathbf{V}_{t}-\sqrt{1-\alpha_{t}}\epsilon^{(t)}_{\theta}(\mathbf{V}_{t})}{\sqrt{\alpha_{t}}}\right)+\sqrt{1-\alpha_{t-1}-\sigma^{2}_{t}}\epsilon^{(t)}_{\theta}(\mathbf{V}_{t})+\sigma_{t}\epsilon,\quad t\in\left[1,t_{0}\right].$ (5)
-Here we use DDIM as sampling scheduler song2020DDIM and $\epsilon_{\theta}$ denotes the video model for noise prediction.
-$\sigma_{t}$ determines whether the denoising process is deterministic or probabilistic, we set $\sigma=1$ to encourage diversity of generation results.
+Here $\epsilon_{\theta}$ denotes the video model for noise prediction and $\sigma_{t}$ determines whether the denoising process is deterministic or probabilistic, we set $\sigma=1$ to encourage diversity of generation results.
 Trade-off Between Fidelity and Diversity
-Leveraging noise prior guidance in diffusion model could lead to trade-off problem between generation’s fidelity and diversity meng2021sdedit ; kim2022diffusionclip ; hou2024high , where results that hold more faithfulness towards guidance tend to decline in generation quality or diversity.
-In this task, similar circumstance also occurs as model is required to receive guidance from a series of imperfect renderings while generate a reasonable and dynamic video.
+Leveraging noise prior guidance in diffusion model could lead to a trade-off between generation’s fidelity and diversity (Meng et al., 2021; Hou et al., 2024), where results that hold more faithfulness towards guidance tend to decline in generation quality.
+In this task, similar circumstance also occurs as model is required to be guided by some imperfect renderings while generate a reasonable and dynamic video.
 The key factor to balance the trade-off problem lies in the choice of $t_{0}$.
@@
 Instead, smaller $t_{0}$ leads to well-generated video, but is less aligned with desired camera motion.
-In our experiments, we find larger $t_{0}$ works better for motions with moderate intensity, and for those with relatively drastic move, smaller $t_{0}$ shows preferable performance.
+Empirically, we find larger $t_{0}$ works better for motions with moderate intensity, and for those with relatively drastic move, smaller $t_{0}$ shows preferable performance.
+The process of stage II is illustrated in Fig.
+2.
 4 Experiments
+Table 1: Quantitative comparisons. Our method attains comparable performance with finetuned methods in both video generation quality and camera motion alignment.
+Method Video Quality Motion Accuracy
+FVD $\downarrow$ FID $\downarrow$ IS $\uparrow$ CLIP-SIM $\uparrow$ ATE $\downarrow$ RPE-T $\downarrow$ RPE-R $\downarrow$
+SVD 1107.93 68.51 7.21 0.3095 4.23 1.79 0.021
+MotionCtrl+SVD 810.59 69.03 7.17 0.3076 4.19 1.07 0.012
+CameraCtrl+SVD 951.80 67.59 6.82 0.3138 4.22 1.17 0.013
+CamTrol+SVD 778.46 68.06 7.05 0.3110 4.17 1.07 0.010
+Reference - - - - 3.60 0.89 0.008
+Figure 3: Qualitative comparisons with finetuned methods.
+CamTrol’s outputs align well with complex trajectories from reference video, while others fail to perceive subtle camera pose changes.
+We provide videos in the supplementary materials for a clearer comparison.
+Table 2: Computational analysis of inference process, evaluated under unified settings.
+SVD MotionCtrl CameraCtrl CamTrol ($t_{0}=10$)
+Max GPU memory(MB) 11542 31702 26208 11542
+Time (s) pre-process - - - 56
+inference 11 32 42 8
+Figure 4: Generalization comparisons.
+CamTrol can avoid domain collapse that rises from overfitting on certain datasets, adapting to more general scenes (Left, where input does not resemble real scenes), and maintain video’s dynamics while adhering to desired camera movements (Right).
 4.1 Experimental Settings
 Implementation Details
-The major results presented in this paper are grounded on zeng2023PixelDance .
-To ensure a fair comparison with other state-of-the-art methods, we employ SVD as base model for quantitative evaluation.
-Inversion and generation steps among all methods are set to 25.
-Our method doesn’t require any additional training utilizing camera trajectories.
+We compare our method with state-of-the-art works including MotionCtrl (Wang et al., 2023) and CameraCtrl (He et al., 2024).
+To ensure a fair comparison, we employ SVD (Blattmann et al., 2023a) as base model for all methods.
+SVD is originally trained on resolution of $576\times 1024$, but SVD-based CameraCtrl only support $320\times 576$.
+Since changing the original resolution leads to suboptimal generation quality, we use $576\times 1024$ for MotionCtrl and CamTrol, then resize their outputs into $320\times 576$ for calculating metrics.
+For all methods, the number of frames and the decoding size of SVD are set to 14.
+We use 25 steps for both inversion and generation processes.
 Evaluation Details
-We compare CamTrol with state-of-the-art works: AnimateDiff guo2023animatediff , MotionCtrl wang2023motionctrl and CameraCtrl he2024cameractrl , all three methods are finetune on SVD using specific data.
-In quantitative evaluation, FVD unterthiner2018FVD , FID heusel2017FID and IS saito2020IS are used to assess video generation quality, while CLIPSIM wu2021CLIPSIM quantifies the similarity between generated video and input prompt.
-With regard to the accuracy of camera motion, we adopt ParticleSFMzhao2022particlesfm and produce estimated camera trajectories from generated videos, with the use of Absolute Trajectory Error(ATE) measuring their differences compared to ground truth.
+In quantitative evaluation, FVD (Unterthiner et al., 2018), FID (Heusel et al., 2017) and IS (Saito et al., 2020) are used to assess video generation quality, and CLIPSIM (Wu et al., 2021) quantifies the similarity between generated video and input prompt.
+For the accuracy of camera motion, we adopt ParticleSFM(Zhao et al., 2022) and produce estimated camera trajectories from generated videos, with the use of Absolute Trajectory Error(ATE) measuring their differences compared to ground truth.
 Relative Pose Error (RPE) is calculated to assess between consecutive frames how well the relative motions match expected ones including their transition (RPE-T) and rotation part (RPE-R).
-Settings of evaluation dataset follow those established in MotionCtrl wang2023motionctrl , but include more complex trajectories.
-Specifically, we extract 51 distinct trajectories from RealEstate10krealestate10k , each paired with 10 prompts mentioned in MotionCtrl wang2023motionctrl , resulting in 510 samples in total for assessment.
-As AnimateDiff guo2023animatediff lacks the ability to handle complicated trajectories, it is only included in qualitative analysis part evaluating with 8 basic camera motions.
-Quantitative comparisons contain AnimateDiff guo2023animatediff can be found in Appendix A.3.
+Specifically, we randomly sample 500 prompt-trajectory pairs from RealEstate10k (Zhou et al., 2018), and use the corresponding videos as reference for calculating FVD and FID.
+Since SVD is an image-to-video model, we generate the first frames using Stable Diffusion (Rombach et al., 2022) by text prompts.
+We also compare the results produces by vanilla SVD as a reference.
+For camera motion accuracy, we provide the evaluations on ground truth videos as a lower bound of these metrics.
 4.2 Comparisons with State-of-the-art Methods
-Quantitative Evaluation
-Quantitative evaluations are shown in Table  1.
-Building upon the same base model as MotionCtrl wang2023motionctrl and CameraCtrl he2024cameractrl , our method demonstrates superior performance across all quantitative metrics concerning both video quality and camera motion accuracy.
-Despite being training-free, CamTrol outperforms those that rely heavily on extensive finetuning and large-scale annotated data, attaining the lowest score in ATE, RPE-T and RPE-R.
-Benefit from explicit camera movements modeling and motion inversion, our method is capable to produce videos align best with those tricky trajectories while maintaining their visual qualities.
-Table 1: Quantitative comparisons.
-Method Video Quality Motion Accuracy
-FVD $\downarrow$ FID $\downarrow$ IS $\uparrow$ CLIP-SIM $\uparrow$ ATE $\downarrow$ RPE-T $\downarrow$ RPE-R $\downarrow$
-MotionCtrl wang2023motionctrl 3576.40 239.10 7.58 0.2933 4.006 1.086 0.106
-CameraCtrl he2024cameractrl 2922.99 243.98 8.07 0.2915 4.200 1.487 0.080
-SVD+CamTrol 2832.59 227.36 8.07 0.3100 3.917 0.947 0.017
-Figure 3: Qualitative comparisons with finetuned methods. CamTrol generates videos that adhere to desired camera movements, without compromising their dynamics and rationality.
-Qualitative Analysis
-Qualitative comparisons of different methods are illustrated in Fig.
-3.
-All methods generate videos align well with desired camera motion, however, the constraints imposed by camera trajectories are sometimes too strong, that output videos have to compromise their quality and rationality to satisfy these movements.
-Some examples that demonstrate this are results of Zoom Out and Roll Anticlockwise, which looks more like flat images being zoomed or rolled on 2D plane instead of videos recorded in 3D spaces.
-The reason might lies on the limitations of proprietary dataset used for finetuning realestate10k ; yu2023mvimgnet , where most scenes are static, making pretrained model forget about knowledge of dynamic contents.
-In contrast, CamTrol makes full use of the powerful dynamics in base model and generate plausible videos with camera perspective change utilizing layout prior guidance, handles these circumstances gracefully.
-More comparisons at Appendix A.3.
-4.3 Ablation Study
 Figure 5: Comparison with base model.
@@
 Figure 6: Effectiveness of layout prior. Layout prior guidance compensates for the vacancies (Left) and static contents (Right) caused by point cloud rendering.
+Quantitative Evaluation
+Quantitative evaluations are shown in Table  1.
+In the table, the best results are in bold, and the second best is underlined.
+The performance of vanilla SVD (without motion control) is indicated as SVD, while the lower bound of motion metrics provided by ground truth videos is indicated as Reference.
+In terms of video quality, CamTrol attains comparable performance with methods finetuned on the RealEstate10k dataset.
+For motion accuracy, CamTrol also gets lowset score in ATE and RPE-T/R.
+The quantitative evaluation demonstrates CamTrol’s ability to generate videos with both accurate camera motion and high visual quality.
+Qualitative Analysis
+Qualitative comparisons are illustrated in Fig.
+3.
+The reference trajectory includes zoom, pan and roll.
+While MotionCtrl and CameraCtrl fail to perceive subtle camera motions, resulting in simple pan movement, CamTrol is able to follow the complex trajectory and generate video with correct move.
+We also evaluate the generalization ability of different methods in generating more general scenes and dynamic content, the results are shown in Fig.
+4.
+Since both MotionCtrl and CameraCtrl are finetuned on limited scenes (i.e., real estate videos) with static content, they can hardly generalize to other scenes, such as videos in non-realistic style or videos that include people.
+As illustrate in Fig.
+4, their motion controls in both situations are misaligned with input movements.
+Furthermore, finetuning on such datasets leads to the loss of dynamics, which is a crucial element in video generation.
+In comparison, CamTrol preserves most of the prior knowledge from video base models, enabling it to handle general scenes as well as generate dynamic content.
+Relevant videos are available in the supplementary materials.
+Computational Analysis
+We provide the computational analysis in Table 2, including the maximum GPU memory required and the time consumption for all methods during inference process.
+The evaluations are conducted under unified settings.
+We test on $576\times 320$, with 25 generation steps, the number of frames and decoding size of SVD are set to 14.
+As a training-free method, CamTrol requires no extra GPU memory during inference process compared with base model.
+This makes it save 10-20GB of GPU memory than other methods under the same circumstances, and makes it applicable running on a single RTX 3090.
+The major consumption of CamTrol comes from the time rendering multi-view images.
+Since this process is currently sequential in time, i.e., $t_{0}\to t_{1}\to t_{2}$, a more parallel approach may be more time-efficient.
+We will leave this to future works.
+The results of $576\times 1024$ resolution and more detailed settings can be found in the Appendix.
+4.3 Ablation Study
+Table 3: Quantitative effect of $t_{0}$.
+$t_{0}$ Video Quality Motion Accuracy
+FVD $\downarrow$ FID $\downarrow$ IS $\uparrow$ CLIP-SIM $\uparrow$ ATE $\downarrow$ RPE-T $\downarrow$ RPE-R $\downarrow$
+$t_{0}=5$ 1079.88 68.52 7.14 0.3100 4.17 1.09 0.012
+$t_{0}=10$ 778.46 68.06 7.05 0.3110 4.17 1.07 0.010
+$t_{0}=15$ 754.14 67.98 7.00 0.3107 4.13 1.02 0.008
+Figure 7: Effect of $t_{0}$. Smaller $t_{0}$ encourages
+dynamics while larger $t_{0}$ preserves camera
+movements (Pedestal Down).
+Figure 8: 3D rotation videos at different motion scales. CamTrol supports camera movements over various scales.
 Comparison to Base Model
@@
 The results are shown in Fig.
-4.
+5.
 It could be observed that even provided with prompts indicating how camera should move, base model fail to produce correct results.
 Instead, CamTrol  is able to implement designated motion control without any instructions from text prompts.
+In Table 1, the comparison with vanilla SVD also demonstrate CamTrol’s effectiveness.
 Effectiveness of Layout Prior
@@
 In Fig.
-5, we showcase frames before and after noise prior guidance.
+6, we showcase frames before and after noise prior guidance.
 With camera pose changes, there appears regions unfilled in point cloud and causes blank spaces in rendered images(left part); Besides, due to the static nature of point cloud, rendered images remain stationary(right part).
@@
 To investigate its effect on the output, we conduct experiments with various $t_{0}$ values, relevant results are shown in Fig.
-7.
+8.
 As illustrated, videos generated with larger $t_{0}$ tend to conform better to camera motion requirements, but suffer from decrease in dynamics; On the contrary, smaller $t_{0}$ leads to more plausible generations but fails to meet camera’s requirements, as latents at these timesteps carry more randomness.
-Generalization to Diverse Base Models
-Our proposed CamTrol can be seamlessly plugged and played with most pretrained video diffusion models to achieve training-free camera control.
-We present visual results showcasing its applications based on different video base models, including SVD SVD and VideoFusion luo2023videofusion , in Fig.
+We also provide quantitative evaluations of different $t_{0}$ in Table 3.
+As $t_{0}$ increase, CamTrol produces videos alike to static, camera-moving scenes (which have lower FVD as we use RealEstate10k as reference videos) with higher accuracy in motion control.
+Generalization to Diverse Situations
+Our proposed CamTrol can be seamlessly plugged and played under various scenarios, accommodating different base models, resolutions, generating length, all in training-free style.
+We present visual results of its applications under different settings, including CogVideoX (Yang et al., 2024b)(diffusion transformer model, $720\times 480$ resolution, 49 frames) and VideoFusion (Luo et al., 2023)(decomposed diffusion process, $128\times 128$ resolution, 16 frames), in Fig.
+10.
+Our approach remains effective applied to alternative video base models, resolutions and generating lengths, demonstrating its strong robustness and generalization ability.
+4.4 Further Applications
+Figure 9: Multi-trajectory video generation. CamTrol is able to generate videos with different trajectories for one scene.
+Figure 10: Applied onto CogVideoX (Yang et al., 2024b) and VideoFusion (Luo et al., 2023).
+CamTrol can be plugged and played under various situations, accommodating different base models, different resolutions, different generating lengths, all in a training-free manner.
+3D Rotation Videos
+One of the most advantages of our method is it can generate videos with rotating moves and produce outputs similar to 3D generation models (Voleti et al., 2024; Melas-Kyriazi et al., 2024).
+While these 3D models need large-scale training on 3D dataset and could only handle inputs in specific styles, our approach is able to deal with any type of images and achieve this in a completely zero-shot manner.
+One example of this is shown in Fig.
 8.
-Our approach remains effective applied to alternative video base models, demonstrating its strong robustness and generalization ability.
-More results with other base models can be found in Appendix A.5.
-4.4 Further Applications
-Figure 6: Effect of $t_{0}$. Smaller $t_{0}$ encourages
-dynamics while larger $t_{0}$ preserves camera
-movements (Pedestal Down).
-Figure 7: Camera control at different scales. CamTrol supports camera movements over various scales.
-Figure 8: Applied onto SVD SVD and VideoFusion luo2023videofusion . CamTrol could be plugged and played with most video diffusion models to offer camera movement control.
-3D Rotation Videos
-One of the most advantages of our method is it can generate videos rotating around some objects and produce outputs similar to 3D generation models sv3d ; im3d .
-While these 3D models need large-scale training on 3D dataset and could only handle inputs in specific styles, our approach is able to deal with any type of images and achieve this in a completely zero-shot manner.
-Some results are shown in 1.
-We offer more results on multi-view synthesis in Appendix A.2.
 Hybrid and Complex Camera Movements
 By combining different basic camera trajectories, CamTrol can support hybrid camera movements and endow generated video with cinematic charm.
-Besides this, explicit motion modeling also equips CamTrol with the abilities to support trajectories containing precise coordinates, which means it can generate videos presenting any complicated camera movement.
-Results about hybrid and complex motions are shown in Fig.
-1 and Appendix A.2.
+Besides this, explicit motion modeling also equips CamTrol with the ability to support trajectories containing precise extrinsics, and generate videos presenting any complicated camera movement.
+Videos avaiable at https://lifedecoder.github.io/CamTrol/.
+Multi-trajectory Video Generation
+One natural application of our method is to generate multi-trajectory videos for one scene.
+Since the point cloud is fixed once the reconstruction process finished, the consistency of the content between different trajectories is guaranteed, and we can easily produce multiple camera-moving videos of this scene.
+We present this application in Fig.
+9.
+More results on multi-trajectory video generation could be found in the supplementary materials.
 Camera Motion at Different Scales
 CamTrol supports camera movements at controllable scales.
-We provide some results in Fig.
-7.
 By specifying different magnitudes of camera’s extrinsic matrix within point cloud spaces, rendered images will exhibit varying degrees of motion, leading to videos with distinct scales of camera movements.
-This further demonstrate the powerful controllability of CamTrol, and provides a new pathway for video’s customized camera control.
+This further demonstrate the powerful controllablity of CamTrol, and provides a new pathway for video’s customized camera control.
+We provide relevant visualization in Fig.
+8.
 5 Conclusion
@@
 It consists of two-stage procedure including explicit camera motion modeling in 3D point cloud space and video generation utilizing layout prior of noisy latents.
-Compared to previours work, CamTrol  does not require any additional finetuning on camera-annotated datasets, or self-supervised training via data augmentation, instead, it could be plugged and played with most video diffusion models to generate camera controllable videos with single image or text prompt as input.
-Comprehensive experimental results demonstrate the effectiveness and robustness of CamTrol for controlling camera motion for videos.
-Besides, we show that CamTrol produces impressive results in generating videos with 3D rotations, complicated trajectories, and different moving scales.
-Appendix A Appendix
-A.1 Definitions of Basic Camera Motions
+Compared to previous work, CamTrol  does not require any additional finetuning on camera-annotated datasets or self-supervised training via data augmentation, simultaneously, comprehensive experiments demonstrate its superior performance in both generation video quality and motion alignment against other state-of-the-art methods.
+Furthermore, we show the ability of CamTrol  generalizing to various scenarios, as well as its impressive applications including unsupervised generation of 3D rotation video, scalable motion control and dealing with complicated camera trajectories.
+Appendix A More Results on Camera Control
+We showcase additional qualitative results of CamTrol on various camera movements. As pdf format is not the best way to exhibit videos, we put them on our demo page:
+If not accessible, please refer to the supplementary materials for a quicker view to the demo page.
+This demo page includes CamTrol-generated videos including basic camera motions (including Zoom, Tilt, Pan, Pedestal, Truck, Roll, Rotate, Hybrid, Complicated, detailed definitions are in B), hybrid motions (Zoom In first, then Pedestal Up, Zoom Out + Pedestal Up + Truck Left + Tilt Down + Pan Right) and complicated motions generated from precise camera extrinsics (extracted from RealEstate10k (Zhou et al., 2018)).
+Besides, it contains 3D rotation videos generated unsupervisedly from video base models (both objects and scenes).
+These outputs share sort of similarity with outputs of 3D generation models, as they all exhibit in a turning-table like way, which camera rotates around some objects.
+The difference here is that 3D model, as trained on specific datasets, could only generate outputs in certain styles, e.g., single static object with no background.
+Instead, our model could handle arbitrary image as input, and generate a rotating-around video with proper dynamics.
+From this aspect, our method could be seen as a infinite source of attaining 3D data.
+And by utilizing our method with stronger backbones, video foundation models could truly become the largest source of 3D data as it should be.
+Furthermore, it contains additional results mentioned in ablation studies, e.g.
+controlling camera motions at different scales, and the effectiveness of using layout prior compared with the raw output given by video base model.
+Appendix B Definitions of Basic Camera Motions
 We refer to the terminology in cinematography to describe different camera motions, definitions of each type are detailed in Table 5.
-Table 2: Definitions of basic camera motions.
+To get consistent images from multiple views, we set camera motion as a trajectory of extrinsic matrix $\left\{\mathbf{P}_{1},...,\mathbf{P}_{N-1}\right\}$, each including a rotation matrix and translation matrix representing camera’s pose and position.
+For hybrid motions, CamTrol supports both spatial (i.e.
+moving several basic moves simultaneously) and temporal (concatenation of basic motions sequentially) combinations.
+In the case of complicated trajectories, one can directly use precise parameters for camera’s extrinsic matrix as input to control video’s motion.
+Additionally, if these parameters not available, a reference video with corresponding move can also serve as input.
+With the help of SFM algorithms (e.g.
+COLMAP  [^1] https://github.com/colmap/colmap, ParticleSFM (Zhao et al., 2022)), camera motion can be easily estimated and used for imitation.
+In this sense, CamTrol can also be seen as an unsupervised method to achieve video motion transfer.
+Appendix C Visual Comparisons with State-of-the-arts
+Sec.
+3 showcases some qualitative comparisons between CamTrol and different state-of-the-art approaches.
+For better visualization and comparison, we provide more video results in the supplementary materials.
+Please check them if needed.
+Appendix D Implementation Details
+For text prompt input, we use Stable Diffusion v2-1  [^2] https://huggingface.co/stabilityai/stable-diffusion-2-1 or Stable Diffusion XL  [^3] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0 to generate the initial image.
+The inpainting model we apply is Stable Diffusion inpainting model proposed by runway  [^4] https://huggingface.co/runwayml/stable-diffusion-inpainting, and backward step of inpainting is set to 25.
+We deploy ZeoDepth  [^5] https://github.com/isl-org/ZoeDepth as depth estimation model.
+The classifier-free guidance scale follows default setting of base models themselves.
+We use SVD’s default setting of 6 fps for video generation, and process reference videos to 6 fps for FVD and FID calculation.
+The complete procedure of CamTrol is elucidated in Algorithm 1.
+For computational analysis, we set both number of frames and decoding size to 14, generation steps to 25.
+We do not apply xformers in all approaches.
+The computational analysis at $576\times 1024$ resolution is shown in Table 4.
+SVD-based CameraCtrl only support resolution at $320\times 576$.
+Table 4: Computational analysis of on $\textbf{576}\times\textbf{1024}$.
+SVD MotionCtrl CameraCtrl CamTrol ($t_{0}=10$)
+Max GPU memory(MB) 34236 71096 - 34236
+Time (s) pre-process - - - 149
+inference 32 54 - 22
+Table 5: Definitions of camera motions. We follow the terminology in cinematography to describe different camera movements.
 Camera Motion Directions Definition
@@
 Anticlockwise
-Hybrid Arbitrary Combination of other motions.
-A.2 More Results on Camera Control
-In this section, we showcase additional qualitative results of CamTrol on various camera movements.
-Fig.
-9 shows video frames with camera move Rotate Anticlockwise and Rotate Clockwise.
-These outputs share sort of similarity with outputs of 3D generation models, as they all exhibit in a turning-table like way, which camera rotates around some objects.
-The difference here is that 3D model, as only trained on specific datasets, could only generate outputs in certain styles, e.g., single static object with no background.
-Instead, our model could handle arbitrary image as input, and generate a rotating-around video with proper dynamics.
-From this aspect, our method could be seen as a infinite source of attaining 3D data.
-And by utilizing our method with stronger backbones, video foundation models could truly become the largest source of 3D data as it should be.
-Besides, we provide additional results on hybrid motions and basic motions, which is illustrated in Fig.
-10, Fig.
-11 and Fig.
-13, respectively.
-We also present pre-loaded complex camera motions in Fig.
-12.
-A.3 More Comparisons to State-of-the-Art Methods
-More comparison results with state-of-the-art methods are shown in Fig.
-14 and Fig.
-15.
-Compared to other works, our method matches well with input texts and camera requirements, meanwhile has wide-ranging diversities on generated contents.
-We also present quantitative comparisons that include AnimateDiff guo2023animatediff as part of the evaluation.
-As AnimateDiff guo2023animatediff lacks the ability to handle complicated trajectories, experiments are conducted focusing on 8 basic trajectories described in A.1.
-To expand the diversity of the models evaluated, we employ CamTrol on zeng2023PixelDance for assessment.
-With regard to basic trajectories, we apply UCF-101 soomro2012UCF as the reference of calculating FID and FVD, from where we randomly sample 125 prompts per camera motion type and form 1000 samples in total for assessment.
-Relevant results are exhibited in Table 3.
-Table 3: Quantitative comparisons on basic trajectories.
-Method Video Quality
-FVD $\downarrow$ FID $\downarrow$ IS $\uparrow$ CLIP-SIM $\uparrow$
-AnimateDiff guo2023animatediff 1582.93 73.79 13.39 0.3196
-MotionCtrl wang2023motionctrl 1768.25 78.26 14.60 0.3143
-CameraCtrl he2024cameractrl 1383.68 76.18 14.04 0.3123
-CamTrol+zeng2023PixelDance 1134.71 69.70 15.87 0.3253
-A.4 Ablation on Depth Optimization
-Ablation results relevant to the effectiveness of depth optimization are presented in Table 4.
-While it affects little on CLIPSIM, depth optimization exhibits benefits on other video quality assessments(FVD, FID, IS) and camera trajectory criteria(ATE, RPE).
-Experiments are conducted on complicated 10 trajectories extracted from RealEstate10k realestate10k , each with 10 prompts mentioned in MotionCtrl wang2023motionctrl .
-We assume that depth optimization offers better alignment between adjacent views rendered from point cloud, which benefits both perceptual quality and camera motion accuracy of generated videos.
-Table 4: Ablation on depth coefficient optimization.
-[…]
-w/o optimization 5033.44 219.01 5.67 0.2949 4.093 1.131 0.046
-w/ optimization 4977.88 218.36 5.72 0.2947 3.906 0.978 0.049
-A.5 More Results on other Base Video Models
-Our methods could be plugged and played with arbitrary video diffusion models to achieve camera control.
-We showcase its performance based on two alternative video generation backbones including SVD SVD and VideoFusion luo2023videofusion , relevant results are exhibited in Fig.
-16.
-Figure 9: Results with camera move Rotate Anticlockwise and Rotate Clockwise. Our method could generated 3D rotation-like videos with dynamics.
-Figure 10: Results with camera move Hybrid I. Here we combine Zoom Out, Truck Left, Pedestal Up, Pan Right and Tilt Down together, form a cinematic like effect on generated videos.
-Figure 11: Results with camera move Hybrid II. Here the hybrid motion is defined as Zoom In first, then Pedestal Up.
-Figure 12: Results with complex camera motions.
-Figure 13: Results with basic camera motions.
-Figure 14: Comparison to state-of-the-art methods on Truck Left, Truck Right, Pedestal Up, Pedestal Down.
-Figure 15: Comparison to state-of-the-art methods on Zoom In, Zoom Out, Roll Anticlockwise, Roll Clockwise.
-Figure 16: Results with other video generation base models including SVD SVD and VideoFusion luo2023videofusion . Our method works well with these backbones.
+Hybrid Combinations Spatial and temporal combination of basic motions.
+Complicated Arbitrary Camera extrinsic matrix or a reference video as input.
+Appendix E Choice of 3D Representation
+Input: Text prompt $p$, camera motion $\mathbf{P}$, input image $\mathbf{I}_{0}$ (optional).
+// Stage I: Camera Motion Modeling:
+1 for i=1,…,N-1 do
+2       $\tilde{\mathbf{I}}_{i}=\text{inpainting}(\mathbf{I}_{i-1},\mathbf{P}_{i},p)$ ;
+3       $\tilde{\mathbf{D}}_{i}=\text{depth}(\tilde{\mathbf{I}}_{i})$ ;
+4       while not converged do
+5             $d_{i}=\operatorname*{argmin}_{d}\left(\sum_{M}\left\|\phi([\tilde{\mathbf{I}}_{i},d\tilde{\mathbf{D}}_{i}],\mathbf{K},\mathbf{P}_{i})-\mathcal{P}_{i-1}\right\|\right)$
+6       end while
+7      $\mathcal{P}_{i}=\phi([\tilde{\mathbf{I}}_{i},d_{i}\tilde{\mathbf{D}}_{i}],\mathbf{K},\mathbf{P}_{i})$ ;
+8       $\mathbf{I}_{i}=\psi(\mathcal{P}_{i},\mathbf{K},\mathbf{P}_{i})$ ;
+9
+10 end for
+// Stage II: Layout Prior Generation:
+11 $\mathbf{V}_{0}\leftarrow\left\{\mathbf{I}_{i}\right\}_{i=0}^{N-1}$ ;
+12 Sample random noise $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ ;
+13 Motion inversion $\mathbf{V}_{t_{0}}\leftarrow\sqrt{\bar{\alpha}_{t_{0}}}\mathbf{V}_{0}+\sqrt{1-\bar{\alpha}_{t_{0}}}\epsilon$ ;
+14 for t=$t_{0}$,…,1 do
+15       $\mathbf{V}_{t-1}\leftarrow\sqrt{\alpha_{t-1}}\left(\frac{\mathbf{V}_{t}-\sqrt{1-\alpha_{t}}\epsilon^{(t)}_{\theta}(\mathbf{V}_{t})}{\sqrt{\alpha_{t}}}\right)+\sqrt{1-\alpha_{t-1}-\sigma^{2}_{t}}\epsilon^{(t)}_{\theta}(\mathbf{V}_{t})+\sigma_{t}\epsilon$
+16 end for
+Algorithm 1 Training-free camera control for video generation
+In Sec.
+3.1, we designate point cloud as the intermediate 3D representation for explicit camera motion modeling.
+Doubts may arise why we do not use more complex 3D representation which might be more useful.
+Here we take the most prevalent 3D representation: 3D Gaussian Splatting  [^6] https://repo-sam.inria.fr/fungraph/3d-gaussian-splatting/, as example to elaborate the benefit of using point cloud in three aspects:
+Firstly, concerning the input, point cloud is able to construct the whole scene from one single input image combining techniques of depth estimation and inpainting.
+GS, though also an explicit 3D representation, requires optimization following images from different views, which means it is neither capable of handling single input image, nor can it leverages 2D inpainting to complete the scene gradually.
+Secondly, from the aspect of time, point cloud can directly lift 2D points onto 3D spaces, while 3DGS demands optimization on each scenario.
+As a training-free method, our method takes nearly no time to generate a video after multi-view images are acquired, but would need more time if 3DGS were applied.
+Lastly, from the task itself, the target of stage I is not precisely reconstruct the 3D scene but only offer a rough layout guidance, in this context, rendered images from point cloud qualify enough and no further refinement on 3D reconstruction is necessary.
+From the analysis, point cloud is quite qualified serving as a rough layout guidance in a relatively quick speed, without any further optimization and redundant multi-view images as input.
+The use of point cloud allows our algorithm to be totally training-free and optimization-free, simultaneously being able to produce camera-moving videos quickly with merely one single image or text prompt as input.
+Appendix F Details about Generating 3D Rotation Videos
+3D generation models (Voleti et al., 2024; Melas-Kyriazi et al., 2024; Shi et al., 2023; Chen et al., 2024; Han et al., 2024) are trained on highly-regulated 3D datasets, these datasets are hard to collect, and consist only a limited variety of data types (e.g.
+single static objects with no background).
+As a consequence, 3D generation models exhibit very constrained output distribution and could only produce results in certain styles.
+CamTrol avoid these problems of arduous data collecting, laborious finetuning and output collapse by utilizing the layout prior hidden in video foundation models.
+Not only does it require no training, but this advantage also benefits CamTrol from inheriting most of the prior knowledge inside video foundation models, e.g., the diversity of scenarios, the dynamic of moving objects, temporal consistency, etc.
+Thus CamTrol is able to generate dynamic 3D content, both objects and scenes, in a totally unsupervised and training-free manner, this is what other method cannot achieve yet.
+Compared with regulated datasets, the problem of processing wild pictures in 3D is that some of the parameters is unknown.
+In Sec.
+3.1, we’ve mentioned that the camera intrinsic matrix $\mathbf{K}$ and initial extrinsic matrix $\mathbf{P}_{0}$ are set by convention as they’re usually intractable.
+Another crucial parameter concerning 3D video generation is the distance between camera and the content of input image (denote as $f$), note that input image could be synthetic or real.
+Considering most camera rotations are done around the center point, we extract a patch from the very center of the input image, conducting depth estimations on the it and define the distance $f$ as the averaged depth.
+The rotation and transition matrix during camera moving can be formed as :
+$\mathcal{R}_{y}$ $=\begin{bmatrix}\cos\theta_{i}&0&-\sin\theta_{i}\\ 0&1&0\\ \sin\theta_{i}&0&\cos\theta_{i}\end{bmatrix},\quad t=\begin{bmatrix}f\sin\theta_{i}\\ 0\\ f-f\cos\theta_{i}\end{bmatrix},$ (6)
+$\text{where}\quad f=\frac{1}{|P|}\sum_{(j,k)\in P}D(x_{0}+j,y_{0}+k).$
+Here $i\in[0,N-1]$ and $\theta_{i}$ refers to rotation angle around $y$ axis at step $i$, $D$ denotes depth estimation of image, $P$ represents the patch around central point $(x_{0},y_{0})$.
+In our experiment, we choose $(j,k)\in[-10,10]$ as the size of patch.
+Appendix G Motion Blur, Problems and Solutions
+Videos produced by CamTrol need to satisfy certain camera movement, and some drastic perspective changes might cause visible trails, recognised as motion blur of objects or scenes.
+This phenomenon will appear to be more indispensable when video base model holds a relatively small generation length (e.g.
+16 frames) as well as the motion scale becomes larger (e.g.
+tilt up for 60 degrees or more).
+To avoid blur issues in controlling video camera motion, we propose several solutions as below:
+  1. 1.
+    According to the analysis above, the blur issue is caused by limited generation frames and large camera movement, thus the most intuitive solution is to either cut down motion scale or utilize a more capable generation model.
+For severe perspective changes, the optimal approach would involve employ video foundation models that support larger generation length (e.g.
+CogVideoX (Yang et al., 2024b) supports generating video with 49 frames).
+This allows model to manage motions of equivalent magnitude while experiencing a smaller moving range between adjacent frames, thereby brings effective alleviation to blur problems.
+  2. 2.
+    One can also stack the results of multiple generations to form a complete outcome, i.e., treating the last frame of previous generation as the starting frame for the next, and integrate them as a whole.
+This approach is more suitable when utilizing a image-to-video (I2V) base model.
+Since most open-source video foundation models are text-to-video (T2V), one may consider increasing the step of camera motion inversion $t_{0}$, which guarantees more fidelity towards input images’ content (and motion).
+  3. 3.
+    Besides the above two approaches, it is as well a common and convenient choice to apply frame interpolation towards the output.
+Lots of off-the-shelf frame interpolation models are open-sourced and could be found on github.
+  4. 4.
+    Lastly, if video length is not being able to alter, it may be necessary to increase the fps of generated videos for better visualization.
+Although larger fps leads to shorter video duration, it simultaneously makes the visual persistence brought by motions less pronounced, which reduce blur visually.
+In our experiments, we take raw output in all settings.

### v4

@@
 Unlike previous work, our method does not require any supervised finetuning on camera-annotated datasets or self-supervised training via data augmentation.
-Instead, it can be plugged and played with most pretrained video diffusion models and generate camera controllable videos with a single image or text prompt as input.
-The inspiration of our work comes from the layout prior that intermediate latents hold towards generated results, thus rearranging noisy pixels in them will make output content reallocated as well.
-As camera move could also be seen as a kind of pixel rearrangement caused by perspective change, videos could be reorganized following specific camera motion if their noisy latents change accordingly.
-Established on this, we propose our method CamTrol, which enables robust camera control for video diffusion models.
+Instead, it can be plug-and-play with most pretrained video diffusion models and generate camera-controllable videos with a single image or text prompt as input.
+The inspiration for our work comes from the layout prior that intermediate latents encode for the generated results, thus rearranging noisy pixels in them will cause the output content to relocate as well.
+As camera moving could also be seen as a type of pixel rearrangement caused by perspective change, videos can be reorganized following specific camera motion if their noisy latents change accordingly.
+Building on this, we propose CamTrol, which enables robust camera control for video diffusion models.
 It is achieved by a two-stage process.
 First, we model image layout rearrangement through explicit camera movement in 3D point cloud space.
-Second, we generate videos with camera motion using layout prior of noisy latents formed by a series of rearranged images.
-Extensive experiments have demonstrated its superior performance in both video generation quality and camera motion alignment compared with other finetuned methods.
-Furthermore, we show the capacity of CamTrol generalizing to various base models, as well as its impressive applications in scalable motion control, dealing with complicated trajectories and unsupervised 3D video generation.
+Second, we generate videos with camera motion by leveraging the layout prior of noisy latents formed by a series of rearranged images.
+Extensive experiments have demonstrated its superior performance in both video generation and camera motion alignment compared with other finetuned methods.
+Furthermore, we show the capability of CamTrol to generalize to various base models, as well as its impressive applications in scalable motion control, dealing with complicated trajectories and unsupervised 3D video generation.
 Videos available at https://lifedecoder.github.io/CamTrol/.
@@
 As a more appealing and content-richer modality, videos differ from images by including an extra temporal dimension.
-This temporal aspect provides increased versatility for depicting diverse and dynamic movements, which can be decomposed into object motion, background transitions and perspective changes.
-Recent years have witnessed the rapid development and splendid breakthrough of video generation with text prompt or images as input instructions (Li et al., 2023; Hong et al., 2022; Ho et al., 2022; Luo et al., 2023; Zeng et al., 2024; Blattmann et al., 2023a; Brooks et al., 2024; Ge et al., 2023; Fei et al., 2023), and demonstrated the inestimable potential of diffusion models to synthesis realistic videos.
+This temporal aspect provides increased versatility for depicting diverse and dynamic movements, which can be decomposed into object motion, background transitions, and perspective changes.
+Recent years have witnessed the rapid development and splendid breakthroughs of video generation with text prompts or images as input instructions (Li et al., 2023; Hong et al., 2022; Ho et al., 2022; Luo et al., 2023; Zeng et al., 2024; Blattmann et al., 2023a; Brooks et al., 2024; Ge et al., 2023; Fei et al., 2023), demonstrating the immense potential of diffusion models to synthesize realistic videos.
 While these video generation models have made progress in generating videos with highly dynamic objects and backgrounds (Zeng et al., 2024; Blattmann et al., 2023a; Li et al., 2023), most of them fail to provide camera control for the generated videos.
-The difficulty of controlling camera trajectory in videos primarily arises from two aspects.
-The initial challenge lies in the inadequacy of annotated data.
-Most video annotations lack of descriptions, especially precise descriptions of video’s camera movements.
+The difficulty of controlling camera motion in videos arises primarily from two aspects.
+The initial challenge lies in the inadequacy of annotated data, as most video annotations lack detailed descriptions, particularly about the camera movements.
 As a result, video generation models trained on these data often fail to interpret text prompts related to camera motions and generate correct outputs.
 One solution to mitigate the data insufficiency problem is to mimic videos with camera movements through simple data augmentation (Yang et al., 2024a).
-However, these methods could only handle simple camera motions like zoom or truck, and have trouble in dealing with more complicated ones.
-The second challenge lies in the effort of additional finetuning required for controlling camera movements.
+However, these methods can only handle simple camera motions like zoom or truck, and have trouble dealing with more complicated ones.
+The second challenge is the extra finetuning required for camera control and its inherent limitations.
 As camera trajectories could be sophisticated, they sometimes cannot be accurately elaborated using naïve text prompts alone.
 Common solutions (Wang et al., 2023; He et al., 2024) proposed to embed camera parameters into diffusion models through learnable encoders and perform extensive finetuning on large-scale datasets with detailed camera trajectories.
-However, such datasets like RealEstate10k (Zhou et al., 2018) and MVImageNet (Yu et al., 2023) are intensively limited in scale and diversity due to the difficulty associated with data collection, in this way, these finetuning methods demand substantial resources but exhibit limited generalizability to other types of data.
-Lack of annotations and heavy finetuning effort make camera control a challenging task in video generations.
+However, such datasets as RealEstate10k (Zhou et al., 2018) and MVImageNet (Yu et al., 2023) are highly limited in scale and diversity due to the difficulty associated with data collection; in this way, these finetuning methods demand substantial resources but exhibit limited generalizability to other types of scenes.
+Lack of annotations and the constraints of finetuning make camera control a challenging task in video generations.
 In this work, we attempt to address these issues through a training-free solution to offer camera control for off-the-shelf video diffusion models.
-We begin by introducing two core observations underpin that video diffusion models can achieve camera movement control in a training-free manner.
+We begin by introducing two core observations underpinning the idea that video diffusion models can achieve camera movement control in a training-free manner.
 First, we find that base video models could produce results with rough camera moves by integrating specific camera-related text into input prompts, such as camera zooms in or camera pans right.
-This simple implementation, though not very accurate and always leads to static or wrong motions, shows the natural prior knowledge learnt by pretrained models about following different camera trajectories.
-The other observation is the effectiveness video models have exhibited in adapting to 3D generation tasks.
-Recent works (Voleti et al., 2024; Melas-Kyriazi et al., 2024; Shi et al., 2023) find that leveraging pretrained video models as initialization helps drastically improve the performance of multi-view generations, demonstrating their strong ability of handling perspective change.
-The two crucial observations reveals the hidden power of video models for camera motion control, thus, we seek to find a way to evoke this innate ability, as it already exists in the model itself.
+This simple implementation, though not very accurate and always leading to static or wrong motions, shows the natural prior knowledge learned by pretrained models about following different camera trajectories.
+The other observation is the effectiveness of video models in adapting to 3D generation tasks.
+Recent works (Voleti et al., 2024; Melas-Kyriazi et al., 2024; Shi et al., 2023) find that leveraging pretrained video models as initialization helps drastically improve the performance of multi-view generations, demonstrating their strong ability to handle perspective change.
+The two crucial observations reveal the hidden power of video models for camera motion control.
+Therefore, we seek to find a way to invoke this innate ability, as it already exists in the model itself.
 We propose CamTrol, which offers camera control for off-the-shelf video diffusion models in a training-free but robust manner.
-CamTrol is inspired by the layout prior that noisy latents hold towards generation results: As pixels in noisy latents change their positions, corresponding rearrangement will also occur to the output and leads to layout modification.
-Considering camera moves could also be seen as a kind of layout rearrangement, this prior can serve as an efficacious hint providing video model with information of specific camera motions.
+CamTrol is inspired by the layout prior that noisy latents possess for the generation results: As pixels in noisy latents change their positions, corresponding rearrangement will also occur in the output, leading to layout modification.
+Considering camera moves can also be seen as a type of layout rearrangement, this prior can serve as an effective hint, providing the video model with information about specific camera motions.
 Specifically, CamTrol consists of a two-stage procedure.
-In stage I, explicit camera movements are modeled in 3D point cloud representations and produce a series of rendered images indicating specific camera movements.
-In stage II, layout prior of noisy latents are utilized to guide video generations with camera movements.
-Compared with previous works, CamTrol require no any additional finetuning utilizing camera annotated datasets, nor does it need self-supervised training based on data augmentation.
+In stage I, explicit camera movements are modeled in 3D point cloud representation and produce a series of rendered images indicating specific camera movements.
+In stage II, the layout prior of noisy latents is utilized to guide video generations with camera movements.
+Compared with previous works, CamTrol requires no additional finetuning on camera-annotated datasets, nor does it need self-supervised training based on data augmentation.
 Extensive experiments have demonstrated its superior performance in both video generation quality and camera motion alignment against other finetuned methods.
-Furthermore, we show the capacity of CamTrol generalizing to various base models, as well as its impressive applications in scalable motion control, dealing with complicated trajectories and unsupervised 3D video generation.
+Furthermore, we show CamTrol’s capability to generalize to various base models, as well as its impressive applications in scalable motion control, dealing with complicated trajectories and unsupervised 3D video generation.
 2 Related Work
 Camera Control for Video Generation
-While methods aim for controlling video foundation models constantly emerge (Ma et al., 2024; Liu et al., 2023; Feng et al., 2023), there are few works explore how to manipulate camera motion of generated videos.
-Earlier work (Hao et al., 2018) controls motion trajectory via warping image through densified sparse flow and pixel fusion, similar ideas also appear later in Chen et al.
+While methods aiming to control video foundation models continue to emerge (Ma et al., 2024; Liu et al., 2023; Feng et al., 2023), there are few works exploring how to manipulate the camera motion of generated videos.
+Earlier works (Hao et al., 2018) control motion trajectory via warping images through densified sparse flow and pixel fusion.
+Similar ideas also appear later in Chen et al.
 (2023) and Yin et al.
 (2023).
-Besides utilizing optical flow, two main techniques for implementing video camera control are via self-supervised augmentation or additional finetuning.
+Apart from utilizing optical flow, two main techniques for implementing video camera control are self-supervised augmentation or additional finetuning.
 Yang et al.
-(2024a) disentangles object motion with camera movement and incorporates extra layers to embed camera motions, where model is trained in a self-supervised manner by augmenting input videos to stimulate simple camera movements.
+(2024a) disentangles object motion with camera movement and incorporates extra layers to embed camera motions, where the model is trained in a self-supervised manner by augmenting input videos to simulate simple camera movements.
 He et al.
 (2024) and Wang et al.
-(2023) train an additional camera encoder and integrate the output into temporal attention layers of U-Net.
+(2023) train additional camera encoders and integrate the outputs into the temporal attention layers of U-Net.
 Guo et al.
-(2023) learns new motion pattern via LoRA (Hu et al., 2021) and finetuning with multiple reference videos.
+(2023) learns new motion patterns via LoRA (Hu et al., 2021) and finetuning with multiple reference videos.
 Noise Prior of Latents in Diffusion Model
-One of the most natural advantages of diffusion model comes from its pixel-wise noisy latents formed during denoising process.
-These latents hold strong causuality towards output and directly determine what the result looks like, meanwhile have robust error-resilience as they are perturbed by Gaussian noises across different scales.
-Numerous work have exploited the convenience of this noise prior to attain controllable generation, such as image-to-image translation (Meng et al., 2021), pixel-level manipulation (Nichol et al., 2022), image inpainting (Lugmayr et al., 2022) and semantic editing (Choi et al., 2021; Hou et al., 2024).
-Recent research has shown that even sampled from Gaussian distribution, the initial noise of diffusion process still have significantly influence to the layout of generated contents (Mao et al., 2023).
-In other work, noise prior is used to guarantee temporal consistency among video frames (Luo et al., 2023), or to trade-off between fidelity and diversity of image editing (Kim et al., 2022).
+One of the most natural advantages of diffusion models comes from its pixel-wise noisy latents formed during the denoising process.
+These latents have a strong causal effect on the output and directly determine what the result looks like, meanwhile having robust error resilience as they are perturbed by Gaussian noise across different scales.
+Numerous works have exploited the convenience of this noise prior to attain controllable generation, such as image-to-image translation (Meng et al., 2021), pixel-level manipulation (Nichol et al., 2022), image inpainting (Lugmayr et al., 2022) and semantic editing (Choi et al., 2021; Hou et al., 2024).
+Recent research has shown that, although sampled from a Gaussian distribution, the initial noise of the diffusion process still has a significant influence on the layout of generated content (Mao et al., 2023).
+In other works, noise prior is used to guarantee temporal consistency between video frames (Luo et al., 2023), or to trade off between fidelity and diversity of image editing (Kim et al., 2022).
 Video Model for 3D Generation
-Similar to how most video generation models using the groundwork laid by image foundation models (Blattmann et al., 2023b; Esser et al., 2023; Singer et al., 2022; Wu et al., 2023), training of 3D generation model also relies heavily on pretrained 2D video models (Voleti et al., 2024; Melas-Kyriazi et al., 2024; Shi et al., 2023; Chen et al., 2024; Han et al., 2024).
-These methods either finetunes with rendered videos directly (Blattmann et al., 2023a; Chen et al., 2024; Melas-Kyriazi et al., 2024; Han et al., 2024), or adds camera embedding for each view as extra condition (Voleti et al., 2024; Shi et al., 2023).
-Video foundation models have shown to be particularly beneficial in generating consistent multi-view rendering of 3D objects, demonstrates their inherent abundant prior knowledge for handling camera pose change.
+Similar to how most video generation models use the groundwork laid by image foundation models (Blattmann et al., 2023b; Esser et al., 2023; Singer et al., 2022; Wu et al., 2023), the training of 3D generation models also relies heavily on pretrained 2D video models (Voleti et al., 2024; Melas-Kyriazi et al., 2024; Shi et al., 2023; Chen et al., 2024; Han et al., 2024).
+These methods either finetune with rendered videos directly (Blattmann et al., 2023a; Chen et al., 2024; Melas-Kyriazi et al., 2024; Han et al., 2024), or add camera embedding as an extra condition for each view (Voleti et al., 2024; Shi et al., 2023).
+Video foundation models have been shown to be particularly beneficial in generating consistent multi-view rendering of 3D objects, demonstrating their inherent and abundant prior knowledge for handling camera pose change.
 3 Training-free Camera Control for Video Generation
-Figure 1: Pipeline of CamTrol. In stage I, camera movements are modeled through explicit 3D point cloud . In stage II, layout prior of noisy latents are utilized to guide video generation.
-CamTrol takes two stages to evoke the innate camera control ability hidden in foundation models.
+Figure 1: Pipeline of CamTrol. In stage I, camera movements are modeled through explicit 3D point cloud. In stage II, the layout prior of noisy latents is utilized to guide video generation.
+CamTrol takes two stages to invoke the innate camera control ability hidden in foundation models.
 In Sec.
@@
 In Sec.
-2, we will elaborate on video’s motion control with the guidance of noise layout prior.
+2, we will elaborate on video motion control with the guidance of noise layout prior.
 3.1 Camera Motion Modeling
-To evoke pretrained video diffusion model’s ability of dealing with camera perspective changes, hints of camera motion should be injected to diffusion model in a proper way.
-While simply concatenating camera trajectories with text prompt is incomprehensible for original model, previous works (Wang et al., 2023; He et al., 2024) introduce additional embedder to encode camera parameters and finetune with limited annotated data (Zhou et al., 2018; Yu et al., 2023), which are data-hungry yet lack of generalization ability.
-Other method (Yang et al., 2024a) constructs camera motions by self-supervised augmentations, but could only handle a few easy camera controls.
-Thus, we seek a more efficient and robust way to guide the model towards camera controllable.
-Considering perspective change of video is originally caused by camera movements in 3D space, we resort to 3D representation for providing generation models with explicit motion hints.
-Specifically, we choose point cloud as the intermediate representation, in which space we can expediently manipulate camera poses for simulating diverse movements.
-One extra benefit that point cloud brings is its data-efficiency: By utilizing inpainting techniques, only one single input image is required for the whole point cloud reconstruction, this sidesteps the effort of large-scale finetuning.
+To invoke pretrained video diffusion models’ ability to deal with camera perspective changes, hints of camera motion should be injected into the diffusion model in a proper way.
+While simply concatenating camera trajectories with text prompts is incomprehensible to the original model, previous works (Wang et al., 2023; He et al., 2024) introduce additional embedders to encode camera parameters and finetune with limited annotated data (Zhou et al., 2018; Yu et al., 2023), which are data-hungry yet lack generalization ability.
+Other methods (Yang et al., 2024a) construct camera motions by self-supervised augmentations, but can only handle a few easy camera controls.
+Thus, we seek a more efficient and robust way to guide the model towards being camera-controllable.
+Considering that video perspective changes are originally caused by camera movements in 3D space, we leverage 3D representation to provide generation models with explicit motion hints.
+Specifically, we choose point cloud as the intermediate representation, in which we can easily manipulate camera poses to simulate diverse movements.
+An additional benefit that point cloud brings is its data efficiency: By utilizing inpainting techniques, only one single input image is required for the entire point cloud reconstruction.
+This eliminates the need of large-scale finetuning.
 Point Cloud Initialization
-We start by lifting the pixels in input image plane to 3D point cloud representations.
+We start by lifting pixels in the input image plane into 3D point cloud space.
 In practice, the input image can be either user-defined or created by image generators like Stable Diffusion (Rombach et al., 2022).
-Given an input image $\mathbf{I}_{0}\in\mathbb{R}^{3\times H\times W}$, we first estimate its depth map $\mathbf{D}_{0}$ using off-the-shelf monocular depth estimator ZeoDepth (Bhat et al., 2023).
-By combining image and its depth map, point cloud $\mathcal{P}_{0}$ can be initialized as:
+Given an input image $\mathbf{I}_{0}\in\mathbb{R}^{3\times H\times W}$, we first estimate its depth map $\mathbf{D}_{0}$ using the off-the-shelf monocular depth estimator ZeoDepth (Bhat et al., 2023).
+By combining the image and its depth map, we initialize the point cloud $\mathcal{P}_{0}$ as:
 $\mathcal{P}_{0}=\phi([\mathbf{I}_{0},\mathbf{D}_{0}],\mathbf{K},\mathbf{P}_{0}),$ (1)
@@
 Camera Trajectories
-To get consistent images from multiple views, we set camera motion as a trajectory of extrinsic matrix $\left\{\mathbf{P}_{1},...,\mathbf{P}_{N-1}\right\}$, each including a rotation matrix and translation matrix representing camera’s pose and position.
-At each step $i$, we project the point cloud back to camera plane using function $\psi$ and get a rendered image with perspective change: $\mathbf{I}_{i}=\psi(\mathcal{P}_{i},\mathbf{K},\mathbf{P}_{i})$.
-By calculating extrinsic matrices of corresponding movement, we obtain a series of camera motions including zoom, tilt, pan, pedestal, truck, roll and rotate, enabling flexible camera movements.
-Detailed definitions of these movements are elaborated in Appendix B.
-By combining basic trajectories, hybrid camera movements can be attained and produce videos with cinematic charm.
-What’s more, benefit from explicit camera motion modeling, our method could support trajectories with precise extrinsics, which means it can generate videos with any complicated camera motion.
+To get consistent images from multiple views, we model the camera motion as a trajectory of extrinsic matrices $\left\{\mathbf{P}_{1},...,\mathbf{P}_{N-1}\right\}$, each including a rotation matrix and a translation matrix representing camera pose and position.
+At each step $i$, we project the point cloud back to the camera plane using $\psi$ and get a rendered image with perspective change: $\mathbf{I}_{i}=\psi(\mathcal{P}_{i},\mathbf{K},\mathbf{P}_{i})$.
+By computing the extrinsic matrices of different movements, we obtain a series of camera motions including zoom, tilt, pan, pedestal, truck, roll, and rotation, enabling flexible camera movements.
+Detailed definitions of these movements are provided in Appendix B.
+Basic trajectories are combined to produce hybrid camera movements, adding cinematic appeal to the generated videos.
+Furthermore, benefiting from explicit camera motion modeling, our method can support trajectories with precise extrinsics, which means it can generate videos with arbitrarily complex camera motions.
 Multi-view Consistency
-When perspective changes, there can be vacancies appear as some areas are unoccupied within the point cloud.
-To get more reasonable results, we employ image inpainting model (Rombach et al., 2022) to fill up the holes for new renderings, with a mask distinguishing the known points from nonexistent ones.
-This operation guarantees coherence between known views and novel views in 2D space.
-After inpainting, image is lifted again onto 3D space and gradually complete the whole point cloud representation.
-During this process, points between adjacent views may become misaligned since depth estimator only estimates relative depth, further leads to inconsistency in both 3D point cloud and rendered images.
+When perspective changes, holes can appear as some areas are unoccupied within the point cloud.
+To obtain more reasonable results, we employ image inpainting (Rombach et al., 2022) to fill the gaps in new renderings, with a mask distinguishing the known points from the nonexistent ones.
+This operation ensures coherence between the known and novel views in 2D space.
+After inpainting, the image is lifted back into 3D space and gradually completes the whole point cloud.
+During this process, points between adjacent views may become misaligned since the depth estimator only estimates relative depth, leading to inconsistencies in 3D point cloud and rendered images.
 To avoid this situation, we adopt depth coefficient optimization (Chung et al., 2023) at each step of the camera movement, formed as:
 $d_{i}=\operatorname*{argmin}_{d}\left(\sum_{M}\left\|\phi([\tilde{\mathbf{I}}_{i},d\tilde{\mathbf{D}}_{i}],\mathbf{K},\mathbf{P}_{i})-\mathcal{P}_{i-1}\right\|\right),$ (2)
-where $\tilde{\mathbf{I}}_{i}$ and $\tilde{\mathbf{D}}_{i}$ refer to the inpainted image and its depth map respectively, $d_{i}$ denotes depth coefficient to be optimized, and $M$ refers to the overlapping region between $\mathcal{P}_{i}$ and $\mathcal{P}_{i-1}$, as other areas are not shared for calculating $\ell_{1}$ loss.
-Thus, we get a set of images refer to the input and indicate specific camera movement:
+where $\tilde{\mathbf{I}}_{i}$ and $\tilde{\mathbf{D}}_{i}$ refer to the inpainted image and the corresponding depth map, respectively, $d_{i}$ denotes the depth coefficient to be optimized, and $M$ refers to the overlapping region between $\mathcal{P}_{i}$ and $\mathcal{P}_{i-1}$, as other areas are not shared for calculating $\ell_{1}$ loss.
+Thus, we get a set of images that correspond to the input and indicate specific camera movements:
 $\left\{\mathbf{I}_{0},...,\mathbf{I}_{N-1}\right\}=\left\{\psi(\mathcal{P}_{i},\mathbf{K},\mathbf{P}_{i})|i\in\left[0,N-1\right]\right\}.$ (3)
@@
 Figure 2: Transition of samples between two distinct distributions.
-As the layout-arranged images being inverted by adding random noise, the distribution of their noisy latents will gradually converge to their video counterpart (green area), finally forming nearly Gaussian latents with specific layout change.
-This information is then inherited during video generation process.
-Steps of camera motion inversion determine the trade-off between video diversity and motion fidelity.
+As the layout-arranged images are inverted by adding random noise, the distribution of their noisy latents will gradually converge to that of their video counterpart (green area), eventually forming a nearly Gaussian latent with specific layout change.
+This information is then inherited during the generation process.
+The inversion step determines the trade-off between video diversity and motion fidelity.
 With camera motion modeling, we obtain a sequence $\mathbf{V}_{0}=\left\{\mathbf{I}_{0},...,\mathbf{I}_{N-1}\right\}\in\mathbb{R}^{N\times 3\times H\times W}$ of rendered images adhering to a specific camera trajectory.
-Note that quality of rendered images are not perfect since single input image only leads to sparse point cloud reconstruction, besides, these renderings are static, thus they could not use directly as video frames.
-To form an ideal video, we need to find a way that satisfies three requirements: 1) camera motions should be maintained; 2) video should be encouraged with more dynamics; and 3) quality imperfection should be compensated.
+Note that the quality of rendered images is not perfect since a single input image only leads to sparse point cloud reconstruction.
+Moreover, these renderings are static, thus they cannot be used directly as video frames.
+To form an ideal video, we need to find a way that satisfies three requirements: 1) camera motions should be maintained; 2) the video should be enriched with more dynamics; and 3) quality imperfection should be compensated.
 Camera Motion Inversion
-Recent work on diffusion models have demonstrated the strong controllability of its noisy latents (Meng et al., 2021; Mao et al., 2023), the causality and error-resilience they hold towards final output make them a convenient yet powerful tool for controllable generation of diffusion models.
-Particularly for initial noise, even sampled from Gaussian distribution, it still have significant influence on the layout of generated image (Mao et al., 2023), so that rearranging the noise pixels will make content in output relocate as well.
-For instance, if all pixels in initial noise shift to right by a certain distance, it is likely that generated output reflect a similar shift.
-This reminds us that the impact of camera movement on images could also be regarded as a kind of layout rearrangement, where pixels change their positions caused by viewpoint change.
+Recent work on diffusion models has demonstrated the strong controllability of their noisy latents (Meng et al., 2021; Mao et al., 2023), the causality and error-resilience they possess with respect to the final output make them a convenient yet powerful tool for controllable generation in diffusion models.
+Particularly for initial noise, even when sampled from a Gaussian distribution, it still has significant influence on the layout of the generated image (Mao et al., 2023).
+For instance, if all pixels in initial noise shift to the right by a certain distance, it is likely that the generated output reflects a similar shift.
+This reminds us that the impact of camera movement on images can also be interpreted as a type of layout rearrangement, where pixels change their positions due to viewpoint change.
 In a similar way, videos can be reorganized following camera motion if their noisy latents change accordingly.
 Inspired by this, we first construct a series of noisy latents indicating specific camera movements.
-It can be intuitively done by employing diffusion’s inversion process on rendered image sequence $\mathbf{V}_{0}$.
-Latent at timestep $t_{0}$ can be calculated as follows, where $\bar{\alpha}_{t}$ are variances of the scheduler:
+It can be done intuitively by employing diffusion’s inversion process on the rendered image sequence $\mathbf{V}_{0}$.
+Latent at timestep $t_{0}$ can be calculated as follows, where $\bar{\alpha}_{t}$ is the variance of the scheduler:
 $\mathbf{V}_{t_{0}}=\sqrt{\bar{\alpha}_{t_{0}}}\mathbf{V}_{0}+\sqrt{1-\bar{\alpha}_{t_{0}}}\epsilon,\quad\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (4)
-Because the rendered images $\mathbf{V}_{0}$ share common pixels in certain regions, their latents also have relevance to each other in a way indicating pixels’ move.
-Moreover, while being perturbed with random noise, blank spaces and flawed regions in $\mathbf{V}_{0}$ can be further filled with randomness, providing video model with more possibilities to generate and correct them.
+Because the rendered images $\mathbf{V}_{0}$ share common pixels in certain regions, their latents are also correlated to each other in a way reflecting pixel movement.
+Moreover, by introducing random noise, blank spaces and flawed regions in $\mathbf{V}_{0}$ can be adaptively completed, providing the video model with more possibilities to generate and correct them.
 Video Generation
-After camera motion inversion, noisy latents presenting camera movements are then passed through the backward process of video diffusion model, utilizing their layout controllability to guide video generation.
-Leveraging prior knowledge of base video model, the generation process also bestows video with rational dynamic information.
-In this way, explicit camera movements are injected into video diffusion model in an appropriate and training-free fashion.
+After camera motion inversion, noisy latents representing camera movements are then passed through the backward process of the video diffusion model, leveraging their layout controllability to guide video generation.
+Using prior knowledge of the base model, the generation process also infuses the video with rational dynamic information.
+In this way, explicit camera movements are injected into the video diffusion model in an appropriate, training-free fashion.
 Starting from noisy motion latents at timestep $t_{0}$, the generation step can be represented as:
 $\hat{\mathbf{V}}_{t-1}=\sqrt{\alpha_{t-1}}\left(\frac{\mathbf{V}_{t}-\sqrt{1-\alpha_{t}}\epsilon^{(t)}_{\theta}(\mathbf{V}_{t})}{\sqrt{\alpha_{t}}}\right)+\sqrt{1-\alpha_{t-1}-\sigma^{2}_{t}}\epsilon^{(t)}_{\theta}(\mathbf{V}_{t})+\sigma_{t}\epsilon,\quad t\in\left[1,t_{0}\right].$ (5)
-Here $\epsilon_{\theta}$ denotes the video model for noise prediction and $\sigma_{t}$ determines whether the denoising process is deterministic or probabilistic, we set $\sigma=1$ to encourage diversity of generation results.
+Here $\epsilon_{\theta}$ denotes the video model for noise prediction and $\sigma_{t}$ determines whether the denoising process is deterministic or probabilistic.
+We set $\sigma=1$ to encourage diversity in generation process.
 Trade-off Between Fidelity and Diversity
-Leveraging noise prior guidance in diffusion model could lead to a trade-off between generation’s fidelity and diversity (Meng et al., 2021; Hou et al., 2024), where results that hold more faithfulness towards guidance tend to decline in generation quality.
-In this task, similar circumstance also occurs as model is required to be guided by some imperfect renderings while generate a reasonable and dynamic video.
-The key factor to balance the trade-off problem lies in the choice of $t_{0}$.
-When larger $t_{0}$ applied, generation bears more resemblance to original guidance $\mathbf{V}_{0}$, yet lacks of rationality and dynamics to be an appealing video.
-Instead, smaller $t_{0}$ leads to well-generated video, but is less aligned with desired camera motion.
-Empirically, we find larger $t_{0}$ works better for motions with moderate intensity, and for those with relatively drastic move, smaller $t_{0}$ shows preferable performance.
+Leveraging noise prior guidance in diffusion models could lead to a trade-off between generation fidelity and diversity (Meng et al., 2021; Hou et al., 2024), where results that are more faithful to the guidance tend to degrade in generation quality.
+In this task, similar circumstances also occur, as the model is required to be guided by some imperfect renderings while generating a reasonable video.
+The key factor in balancing the trade-off problem lies in the choice of $t_{0}$.
+When a smaller $t_{0}$ is applied, the generated output bears more resemblance to original guidance $\mathbf{V}_{0}$, but lacks rationality and dynamics to be an appealing video.
+Instead, a larger $t_{0}$ leads to better-generated videos but makes them less aligned with the desired camera motion.
+Empirically, we find that a smaller $t_{0}$ works better for motions with moderate intensity, and for those with relatively drastic movements, a larger $t_{0}$ shows preferable performance.
 The process of stage II is illustrated in Fig.
@@
 Figure 3: Qualitative comparisons with finetuned methods.
-CamTrol’s outputs align well with complex trajectories from reference video, while others fail to perceive subtle camera pose changes.
-We provide videos in the supplementary materials for a clearer comparison.
+CamTrol’s outputs align well with complex trajectories from the reference video, while others fail to capture subtle changes in camera pose.
+We provide videos in the supplementary materials for clearer comparison.
 Table 2: Computational analysis of inference process, evaluated under unified settings.
-SVD MotionCtrl CameraCtrl CamTrol ($t_{0}=10$)
+SVD MotionCtrl CameraCtrl CamTrol ($t_{0}=15$)
 Max GPU memory(MB) 11542 31702 26208 11542
@@
 Figure 4: Generalization comparisons.
-CamTrol can avoid domain collapse that rises from overfitting on certain datasets, adapting to more general scenes (Left, where input does not resemble real scenes), and maintain video’s dynamics while adhering to desired camera movements (Right).
+CamTrol can avoid domain collapse that arises from overfitting on certain datasets and adapt to more general scenes (Left), meanwhile preserving video’s dynamics while adhering to desired camera movements (Right).
 4.1 Experimental Settings
@@
 To ensure a fair comparison, we employ SVD (Blattmann et al., 2023a) as base model for all methods.
-SVD is originally trained on resolution of $576\times 1024$, but SVD-based CameraCtrl only support $320\times 576$.
-Since changing the original resolution leads to suboptimal generation quality, we use $576\times 1024$ for MotionCtrl and CamTrol, then resize their outputs into $320\times 576$ for calculating metrics.
+SVD is originally trained at a resolution of $576\times 1024$, but SVD-based CameraCtrl only supports $320\times 576$.
+Since changing the original resolution results in suboptimal generation quality, we use $576\times 1024$ for MotionCtrl and CamTrol, then resize the outputs to $320\times 576$ to calculate metrics.
 For all methods, the number of frames and the decoding size of SVD are set to 14.
-We use 25 steps for both inversion and generation processes.
+We use 25 steps for both the inversion and generation processes.
 Evaluation Details
-In quantitative evaluation, FVD (Unterthiner et al., 2018), FID (Heusel et al., 2017) and IS (Saito et al., 2020) are used to assess video generation quality, and CLIPSIM (Wu et al., 2021) quantifies the similarity between generated video and input prompt.
-For the accuracy of camera motion, we adopt ParticleSFM(Zhao et al., 2022) and produce estimated camera trajectories from generated videos, with the use of Absolute Trajectory Error(ATE) measuring their differences compared to ground truth.
-Relative Pose Error (RPE) is calculated to assess between consecutive frames how well the relative motions match expected ones including their transition (RPE-T) and rotation part (RPE-R).
-Specifically, we randomly sample 500 prompt-trajectory pairs from RealEstate10k (Zhou et al., 2018), and use the corresponding videos as reference for calculating FVD and FID.
-Since SVD is an image-to-video model, we generate the first frames using Stable Diffusion (Rombach et al., 2022) by text prompts.
-We also compare the results produces by vanilla SVD as a reference.
+In the quantitative evaluation, FVD (Unterthiner et al., 2018), FID (Heusel et al., 2017) and IS (Saito et al., 2020) are used to assess video generation quality, and CLIPSIM (Wu et al., 2021) quantifies the similarity between the generated video and the input prompt.
+For camera motion accuracy, we adopt ParticleSFM(Zhao et al., 2022) to estimate camera trajectories from generated videos, with the use of Absolute Trajectory Error (ATE) to measure their differences from the ground truth.
+Relative Pose Error (RPE) is calculated to assess how well the relative motions between consecutive frames match the expected ones, including their translation (RPE-T) and rotation component (RPE-R).
+Specifically, we randomly sample 500 prompt-trajectory pairs from RealEstate10k (Zhou et al., 2018), and use them as references for calculating FVD and FID.
+Since SVD is an image-to-video model, we generate the first frames using Stable Diffusion (Rombach et al., 2022) based on text prompts.
+We also provide the results produced by vanilla SVD as a reference.
 For camera motion accuracy, we provide the evaluations on ground truth videos as a lower bound of these metrics.
 4.2 Comparisons with State-of-the-art Methods
-Figure 5: Comparison with base model.
-Controlling camera motion via prompt engineering doesn’t work at most times.
-Instead, CamTrol offers robust control towards video’s camera movement in a training-free manner.
-Figure 6: Effectiveness of layout prior. Layout prior guidance compensates for the vacancies (Left) and static contents (Right) caused by point cloud rendering.
+Figure 5: Comparison with base model. Controlling camera motion via prompt engineering rarely succeeds. Instead, CamTrol offers robust control to video’s camera movements.
+Figure 6: Effectiveness of layout prior. Layout prior guidance compensates for the voids (Left) and static elements (Right) in point cloud.
 Quantitative Evaluation
 Quantitative evaluations are shown in Table  1.
-In the table, the best results are in bold, and the second best is underlined.
-The performance of vanilla SVD (without motion control) is indicated as SVD, while the lower bound of motion metrics provided by ground truth videos is indicated as Reference.
-In terms of video quality, CamTrol attains comparable performance with methods finetuned on the RealEstate10k dataset.
-For motion accuracy, CamTrol also gets lowset score in ATE and RPE-T/R.
+In the table, the best results are in bold, and the second best are underlined.
+The performance of vanilla SVD (without motion control) is denoted as SVD, while the lower bound for motion metrics, provided by ground truth videos, is denoted as Reference.
+In terms of video quality, CamTrol attains performance comparable to methods finetuned on the RealEstate10k dataset.
+For motion accuracy, CamTrol also achieves the lowest score in ATE and RPE-T/R.
 The quantitative evaluation demonstrates CamTrol’s ability to generate videos with both accurate camera motion and high visual quality.
@@
 3.
-The reference trajectory includes zoom, pan and roll.
-While MotionCtrl and CameraCtrl fail to perceive subtle camera motions, resulting in simple pan movement, CamTrol is able to follow the complex trajectory and generate video with correct move.
-We also evaluate the generalization ability of different methods in generating more general scenes and dynamic content, the results are shown in Fig.
+The reference trajectory includes zoom, pan, and roll.
+While MotionCtrl and CameraCtrl fail to capture subtle camera motions, resulting in simple pan movements, CamTrol is able to follow the complex trajectory and generate videos with correct motion.
+We also evaluate the generalizability of different methods in generating more general scenes and dynamic content.
+The results are shown in Fig.
 4.
-Since both MotionCtrl and CameraCtrl are finetuned on limited scenes (i.e., real estate videos) with static content, they can hardly generalize to other scenes, such as videos in non-realistic style or videos that include people.
-As illustrate in Fig.
-4, their motion controls in both situations are misaligned with input movements.
-Furthermore, finetuning on such datasets leads to the loss of dynamics, which is a crucial element in video generation.
+Since both MotionCtrl and CameraCtrl are finetuned on limited scenes (i.e., real estate videos) with static content, they have difficulty generalizing to other scenes, such as videos in non-realistic styles or videos that include people.
+As illustrated in Fig.
+4, their motion controls in both situations are misaligned with the intended movements.
+Furthermore, finetuning on such datasets leads to a loss of dynamics, which is a crucial element in video generation.
 In comparison, CamTrol preserves most of the prior knowledge from video base models, enabling it to handle general scenes as well as generate dynamic content.
@@
 Computational Analysis
-We provide the computational analysis in Table 2, including the maximum GPU memory required and the time consumption for all methods during inference process.
-The evaluations are conducted under unified settings.
-We test on $576\times 320$, with 25 generation steps, the number of frames and decoding size of SVD are set to 14.
-As a training-free method, CamTrol requires no extra GPU memory during inference process compared with base model.
-This makes it save 10-20GB of GPU memory than other methods under the same circumstances, and makes it applicable running on a single RTX 3090.
-The major consumption of CamTrol comes from the time rendering multi-view images.
-Since this process is currently sequential in time, i.e., $t_{0}\to t_{1}\to t_{2}$, a more parallel approach may be more time-efficient.
-We will leave this to future works.
-The results of $576\times 1024$ resolution and more detailed settings can be found in the Appendix.
+We provide the computational analysis in Table 2, including the maximum GPU memory required and the processing time for all methods during inference.
+Evaluations are conducted under unified settings.
+We test at a resolution of $576\times 320$ with 25 generation steps.
+The number of frames and the decoding size of SVD are set to 14.
+As a training-free method, CamTrol requires no extra GPU memory during the inference process compared to the base model.
+This saves 10-20GB of GPU memory compared to other methods under the same circumstances, allowing it to run on a single RTX 3090.
+The major consumption of CamTrol comes from rendering multi-view images.
+Since this process is temporally sequential, i.e., $t_{0}\to t_{1}\to t_{2}$, a more parallelized approach may improve time efficiency.
+The results for $576\times 1024$ resolution and more detailed settings can be found in the Appendix.
 4.3 Ablation Study
@@
 FVD $\downarrow$ FID $\downarrow$ IS $\uparrow$ CLIP-SIM $\uparrow$ ATE $\downarrow$ RPE-T $\downarrow$ RPE-R $\downarrow$
-$t_{0}=5$ 1079.88 68.52 7.14 0.3100 4.17 1.09 0.012
-$t_{0}=10$ 778.46 68.06 7.05 0.3110 4.17 1.07 0.010
-$t_{0}=15$ 754.14 67.98 7.00 0.3107 4.13 1.02 0.008
-Figure 7: Effect of $t_{0}$. Smaller $t_{0}$ encourages
-dynamics while larger $t_{0}$ preserves camera
+$t_{0}=20$ 1079.88 68.52 7.14 0.3100 4.17 1.09 0.012
+$t_{0}=15$ 778.46 68.06 7.05 0.3110 4.17 1.07 0.010
+$t_{0}=10$ 754.14 67.98 7.00 0.3107 4.13 1.02 0.008
+Figure 7: Effect of $t_{0}$. Larger $t_{0}$ encourages
+dynamics while smaller $t_{0}$ preserves camera
 movements (Pedestal Down).
-Figure 8: 3D rotation videos at different motion scales. CamTrol supports camera movements over various scales.
+Figure 8: 3D rotation videos at different motion scales. CamTrol supports camera movements across multiple scales.
 Comparison to Base Model
-To demonstrate that changes of camera motion are attributed to our method rather than the innate capability of video model, we conduct ablation study to assess its effectiveness.
-We add prompts describing certain camera moves (e.g.
-zooms out), letting video model understand by itself.
-The results are shown in Fig.
+To demonstrate that the changes in camera motion are attributed to our method rather than the innate capability of the video model, we conduct an ablation study to evaluate its effectiveness.
+We add prompts describing specific camera movements (e.g.
+zooms out) and let the video model interpret them on its own.
+The results are presented in Fig.
 5.
-It could be observed that even provided with prompts indicating how camera should move, base model fail to produce correct results.
-Instead, CamTrol  is able to implement designated motion control without any instructions from text prompts.
-In Table 1, the comparison with vanilla SVD also demonstrate CamTrol’s effectiveness.
+It can be observed that, even when provided with prompts indicating how the camera should move, the base model fails to produce correct results.
+In contrast, CamTrol is able to implement the designated motion control without any instructions from text prompts.
+In Table 1, the comparisons with vanilla SVD also demonstrate CamTrol’s effectiveness.
 Effectiveness of Layout Prior
-We employ ablation study to validate the effectiveness of layout prior guidance, illustrating its necessity from two aspects: the completeness of vacancies and dynamic of generated video.
+We conduct an ablation study to validate the effectiveness of layout prior guidance, demonstrating its necessity in two aspects: the completeness of missing areas and the dynamics of the generated video.
 In Fig.
-6, we showcase frames before and after noise prior guidance.
-With camera pose changes, there appears regions unfilled in point cloud and causes blank spaces in rendered images(left part); Besides, due to the static nature of point cloud, rendered images remain stationary(right part).
-Noise layout prior could compensate for these flaws, finally produce videos with inpainted vacancies and rationalized dynamics.
+6, we showcase frames before and after applying noise prior guidance.
+As camera pose changes, there are regions left unfilled in the point cloud, leading to blank spaces in rendered images (left part); Additionally, due to the static nature of point cloud, the rendered images remain stationary (right part).
+Layout prior guidance compensates for these flaws, ultimately producing videos with inpainted gaps and natural dynamics.
 Effect of Timestep $t_{0}$
-$t_{0}$ is a crucial factor that influences the trade-off between generated video’s diversity and its faithfulness to camera motions requirements.
-To investigate its effect on the output, we conduct experiments with various $t_{0}$ values, relevant results are shown in Fig.
+$t_{0}$ is a crucial factor that influences the trade-off between generated video’s diversity and its faithfulness to camera motion requirements.
+To investigate its effect on the output, we conduct experiments with various $t_{0}$ values, the relevant results are shown in Fig.
 8.
-As illustrated, videos generated with larger $t_{0}$ tend to conform better to camera motion requirements, but suffer from decrease in dynamics; On the contrary, smaller $t_{0}$ leads to more plausible generations but fails to meet camera’s requirements, as latents at these timesteps carry more randomness.
-We also provide quantitative evaluations of different $t_{0}$ in Table 3.
-As $t_{0}$ increase, CamTrol produces videos alike to static, camera-moving scenes (which have lower FVD as we use RealEstate10k as reference videos) with higher accuracy in motion control.
+As illustrated, videos generated with smaller $t_{0}$ tend to conform better to camera motion requirements, but suffer from decreased dynamics; On the other hand, larger $t_{0}$ leads to more plausible generations but fails to meet camera’s requirements, since the latents at these timesteps carry more randomness.
+We also provide quantitative evaluations for different $t_{0}$ values in Table 3.
+As $t_{0}$ decreases, CamTrol produces videos resembling static, camera-moving scenes (which have lower FVD as we use RealEstate10k as the reference videos) with higher accuracy in motion control.
 Generalization to Diverse Situations
-Our proposed CamTrol can be seamlessly plugged and played under various scenarios, accommodating different base models, resolutions, generating length, all in training-free style.
+Our proposed CamTrol can be seamlessly integrated into various scenarios, accommodating different base models, resolutions, and generation lengths, all in a training-free style.
 We present visual results of its applications under different settings, including CogVideoX (Yang et al., 2024b) (diffusion transformer model, $720\times 480$ resolution, 49 frames) and VideoFusion (Luo et al., 2023) (decomposed diffusion process, $128\times 128$ resolution, 16 frames), in Fig.
 10.
-Our approach remains effective applied to alternative video base models, resolutions and generating lengths, demonstrating its strong robustness and generalization ability.
+Our approach remains effective when applied to alternative video base models, resolutions, and generation lengths, demonstrating its strong robustness and generalizability.
 4.4 Further Applications
-Figure 9: Multi-trajectory video generation. CamTrol is able to generate videos with different trajectories for one scene.
-Figure 10: Applied onto CogVideoX (Yang et al., 2024b) and VideoFusion (Luo et al., 2023).
-CamTrol can be plugged and played under various situations, accommodating different base models, different resolutions, different generating lengths, all in a training-free manner.
+Figure 9: Multi-trajectory video generation. CamTrol is able to generate videos with different camera motions for the same scene.
+Figure 10: Applied to CogVideoX (Yang et al., 2024b) and VideoFusion (Luo et al., 2023).
+CamTrol can be plug-and-play under various situations, accommodating different base models, different resolutions, different generation lengths, all in a training-free manner.
 3D Rotation Videos
-One of the most advantages of our method is it can generate videos with rotating moves and produce outputs similar to 3D generation models (Voleti et al., 2024; Melas-Kyriazi et al., 2024).
-While these 3D models need large-scale training on 3D dataset and could only handle inputs in specific styles, our approach is able to deal with any type of images and achieve this in a completely zero-shot manner.
-One example of this is shown in Fig.
+One of the main advantages of our method is that it can generate videos with rotating movements and produce outputs similar to those of 3D generation models (Voleti et al., 2024; Melas-Kyriazi et al., 2024).
+While these 3D models require large-scale training on 3D datasets and can only handle inputs in specific styles, our approach is capable of dealing with any type of image and achieving this in a completely zero-shot manner.
+An example is shown in Fig.
 8.
 Hybrid and Complex Camera Movements
-By combining different basic camera trajectories, CamTrol can support hybrid camera movements and endow generated video with cinematic charm.
-Besides this, explicit motion modeling also equips CamTrol with the ability to support trajectories containing precise extrinsics, and generate videos presenting any complicated camera movement.
-Videos avaiable at https://lifedecoder.github.io/CamTrol/.
+By combining basic camera trajectories, CamTrol can support hybrid camera movements and produce videos with cinematic appeal.
+In addition, explicit motion modeling equips CamTrol with the ability to handle trajectories of precise extrinsics, allowing it to generate videos with arbitrarily complicated camera movements.
+Relevant results can be found at https://lifedecoder.github.io/CamTrol/.
 Multi-trajectory Video Generation
-One natural application of our method is to generate multi-trajectory videos for one scene.
-Since the point cloud is fixed once the reconstruction process finished, the consistency of the content between different trajectories is guaranteed, and we can easily produce multiple camera-moving videos of this scene.
-We present this application in Fig.
+One natural application of our method is to generate multi-trajectory videos for a single scene.
+Since the point cloud remains fixed once the reconstruction has finished, the spatial consistency between different trajectories is guaranteed, making it easy to produce multiple camera-moving videos of the same scene.
+We demonstrate this application in Fig.
 9.
-More results on multi-trajectory video generation could be found in the supplementary materials.
+More results on multi-trajectory video generation can be found in the supplementary materials.
 Camera Motion at Different Scales
 CamTrol supports camera movements at controllable scales.
-By specifying different magnitudes of camera’s extrinsic matrix within point cloud spaces, rendered images will exhibit varying degrees of motion, leading to videos with distinct scales of camera movements.
-This further demonstrate the powerful controllablity of CamTrol, and provides a new pathway for video’s customized camera control.
+By specifying different magnitudes of the camera’s extrinsic matrix, the rendered images will exhibit varying degrees of motion, resulting in videos with distinct scales of camera movements.
+This further demonstrates the powerful controllability of CamTrol and provides a new pathway for customized camera control of video generation.
 We provide relevant visualization in Fig.
@@
 In this paper, we propose a training-free and robust method CamTrol to offer camera control for off-the-shelf video diffusion models.
-It consists of two-stage procedure including explicit camera motion modeling in 3D point cloud space and video generation utilizing layout prior of noisy latents.
-Compared to previous work, CamTrol  does not require any additional finetuning on camera-annotated datasets or self-supervised training via data augmentation, simultaneously, comprehensive experiments demonstrate its superior performance in both generation video quality and motion alignment against other state-of-the-art methods.
-Furthermore, we show the ability of CamTrol  generalizing to various scenarios, as well as its impressive applications including unsupervised generation of 3D rotation video, scalable motion control and dealing with complicated camera trajectories.
+It consists of a two-stage procedure, including explicit camera motion modeling in 3D point cloud space and video generation utilizing the layout prior of noisy latents.
+Compared to previous works, CamTrol does not require additional finetuning on camera-annotated datasets or self-supervised training via data augmentation.
+Comprehensive experiments demonstrate its superior performance in both generation quality and motion alignment against other state-of-the-art methods.
+Furthermore, we show the ability of CamTrol to generalize to various scenarios, as well as its impressive applications including unsupervised generation of 3D video, scalable motion control, and dealing with complicated camera trajectories.
 Appendix A More Results on Camera Control
-We showcase additional qualitative results of CamTrol on various camera movements. As pdf format is not the best way to exhibit videos, we put them on our demo page:
-If not accessible, please refer to the supplementary materials for a quicker view to the demo page.
-This demo page includes CamTrol-generated videos including basic camera motions (including Zoom, Tilt, Pan, Pedestal, Truck, Roll, Rotate, Hybrid, Complicated, detailed definitions are in B), hybrid motions (Zoom In first, then Pedestal Up, Zoom Out + Pedestal Up + Truck Left + Tilt Down + Pan Right) and complicated motions generated from precise camera extrinsics (extracted from RealEstate10k (Zhou et al., 2018)).
-Besides, it contains 3D rotation videos generated unsupervisedly from video base models (both objects and scenes).
-These outputs share sort of similarity with outputs of 3D generation models, as they all exhibit in a turning-table like way, which camera rotates around some objects.
-The difference here is that 3D model, as trained on specific datasets, could only generate outputs in certain styles, e.g., single static object with no background.
-Instead, our model could handle arbitrary image as input, and generate a rotating-around video with proper dynamics.
-From this aspect, our method could be seen as a infinite source of attaining 3D data.
-And by utilizing our method with stronger backbones, video foundation models could truly become the largest source of 3D data as it should be.
-Furthermore, it contains additional results mentioned in ablation studies, e.g.
-controlling camera motions at different scales, and the effectiveness of using layout prior compared with the raw output given by video base model.
+We present additional qualitative results of CamTrol on our demo page:
+This demo page includes CamTrol-generated videos including basic camera motions (including Zoom, Tilt, Pan, Pedestal, Truck, Roll, Rotate, Hybrid, Complicated, detailed definitions are in Appendix B), hybrid motions (Zoom In first, then Pedestal Up, Zoom Out + Pedestal Up + Truck Left + Tilt Down + Pan Right), and complicated motions generated from precise camera extrinsics (extracted from RealEstate10k (Zhou et al., 2018)).
+In addition, it contains 3D rotation videos generated unsupervised from video base models (both objects and scenes).
+These outputs share similarities with those of 3D generation models, as they all exhibit in a turning-table-like way, with the camera rotating around objects.
+The difference here is that the 3D model, as trained on specific datasets, can only generate outputs in certain styles, e.g., single static object with no background.
+Instead, our model can handle arbitrary images as input and generate rotating videos with appropriate dynamics.
+From this perspective, our method could be seen as an infinite source of 3D data.
+By employing our method with stronger backbones, video foundation models could truly become the largest source of 3D data, as they should be.
+Furthermore, it contains additional results discussed in ablation studies, e.g., controlling camera motions at different scales, and the effectiveness of using layout prior compared with the raw output given by the video base model.
 Appendix B Definitions of Basic Camera Motions
-We refer to the terminology in cinematography to describe different camera motions, definitions of each type are detailed in Table 5.
-To get consistent images from multiple views, we set camera motion as a trajectory of extrinsic matrix $\left\{\mathbf{P}_{1},...,\mathbf{P}_{N-1}\right\}$, each including a rotation matrix and translation matrix representing camera’s pose and position.
-For hybrid motions, CamTrol supports both spatial (i.e.
-moving several basic moves simultaneously) and temporal (concatenation of basic motions sequentially) combinations.
+We refer to the terminology in cinematography to describe different camera motions, the definitions of each type are detailed in Table 5.
+To obtain consistent images from multiple views, we set camera motion as a trajectory of extrinsic matrices $\left\{\mathbf{P}_{1},...,\mathbf{P}_{N-1}\right\}$, each consisting of a rotation matrix and a translation matrix, representing the camera pose and position.
+For hybrid motions, CamTrol supports both spatial (i.e., performing several basic motions simultaneously) and temporal (i.e., concatenating basic motions sequentially) combinations.
 In the case of complicated trajectories, one can directly use precise parameters for camera’s extrinsic matrix as input to control video’s motion.
-Additionally, if these parameters not available, a reference video with corresponding move can also serve as input.
-With the help of SFM algorithms (e.g.
-COLMAP  [^1] https://github.com/colmap/colmap, ParticleSFM (Zhao et al., 2022)), camera motion can be easily estimated and used for imitation.
+Additionally, if these parameters are not available, a reference video with corresponding move can also serve as input.
+With the help of SFM algorithms (e.g., COLMAP  [^1] https://github.com/colmap/colmap, ParticleSFM (Zhao et al., 2022)), camera motion can be easily estimated and used for imitation.
 In this sense, CamTrol can also be seen as an unsupervised method to achieve video motion transfer.
@@
 For better visualization and comparison, we provide more video results in the supplementary materials.
-Please check them if needed.
 Appendix D Implementation Details
 For text prompt input, we use Stable Diffusion v2-1  [^2] https://huggingface.co/stabilityai/stable-diffusion-2-1 or Stable Diffusion XL  [^3] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0 to generate the initial image.
-The inpainting model we apply is Stable Diffusion inpainting model proposed by runway  [^4] https://huggingface.co/runwayml/stable-diffusion-inpainting, and backward step of inpainting is set to 25.
-We deploy ZeoDepth  [^5] https://github.com/isl-org/ZoeDepth as depth estimation model.
-The classifier-free guidance scale follows default setting of base models themselves.
-We use SVD’s default setting of 6 fps for video generation, and process reference videos to 6 fps for FVD and FID calculation.
+The inpainting model we apply is Stable Diffusion inpainting model proposed by Runway  [^4] https://huggingface.co/runwayml/stable-diffusion-inpainting, and the backward step of inpainting is set to 25.
+We use ZeoDepth  [^5] https://github.com/isl-org/ZoeDepth as depth estimation model.
+The classifier-free guidance scale follows the default setting of the base models themselves.
+We use SVD’s default setting of 6 fps for video generation, and process reference videos to 6 fps for FVD and FID calculations.
 The complete procedure of CamTrol is elucidated in Algorithm 1.
-For computational analysis, we set both number of frames and decoding size to 14, generation steps to 25.
-We do not apply xformers in all approaches.
+For computational analysis, we set both the number of frames and the decoding size to 14, and the generation steps to 25.
+We do not apply xformers in any of the approaches.
 The computational analysis at $576\times 1024$ resolution is shown in Table 4.
-SVD-based CameraCtrl only support resolution at $320\times 576$.
+SVD-based CameraCtrl only supports a resolution at $320\times 576$.
 Table 4: Computational analysis of on $\textbf{576}\times\textbf{1024}$.
-SVD MotionCtrl CameraCtrl CamTrol ($t_{0}=10$)
+SVD MotionCtrl CameraCtrl CamTrol ($t_{0}=15$)
 Max GPU memory(MB) 34236 71096 - 34236
@@
 inference 32 54 - 22
-Table 5: Definitions of camera motions. We follow the terminology in cinematography to describe different camera movements.
+Table 5: Definitions of camera motions. We adopt the terminology from cinematography to describe different camera movements.
 Camera Motion Directions Definition
@@
 Right
-Pedestal Up Moving a camera vertically in its entirety.
+Pedestal Up Moving the camera vertically in its entirety.
 Down
-Truck Left Moving a camera horizontally in its entirety.
+Truck Left Moving the camera horizontally in its entirety.
 Right
-Roll Clockwise Rotating a camera in its entirety in a horizontal manner.
+Roll Clockwise Rotating the camera in its entirety in a horizontal manner.
 Anticlockwise
-Rotate Clockwise Moving a camera around a subject.
+Rotate Clockwise Moving the camera around a subject.
 Anticlockwise
-Hybrid Combinations Spatial and temporal combination of basic motions.
-Complicated Arbitrary Camera extrinsic matrix or a reference video as input.
+Hybrid Combinations Spatial and temporal combinations of basic motions.
+Complicated Arbitrary Camera extrinsic matrices or reference videos as input.
 Appendix E Choice of 3D Representation
@@
 3.1, we designate point cloud as the intermediate 3D representation for explicit camera motion modeling.
-Doubts may arise why we do not use more complex 3D representation which might be more useful.
-Here we take the most prevalent 3D representation: 3D Gaussian Splatting  [^6] https://repo-sam.inria.fr/fungraph/3d-gaussian-splatting/, as example to elaborate the benefit of using point cloud in three aspects:
-Firstly, concerning the input, point cloud is able to construct the whole scene from one single input image combining techniques of depth estimation and inpainting.
-GS, though also an explicit 3D representation, requires optimization following images from different views, which means it is neither capable of handling single input image, nor can it leverages 2D inpainting to complete the scene gradually.
-Secondly, from the aspect of time, point cloud can directly lift 2D points onto 3D spaces, while 3DGS demands optimization on each scenario.
-As a training-free method, our method takes nearly no time to generate a video after multi-view images are acquired, but would need more time if 3DGS were applied.
-Lastly, from the task itself, the target of stage I is not precisely reconstruct the 3D scene but only offer a rough layout guidance, in this context, rendered images from point cloud qualify enough and no further refinement on 3D reconstruction is necessary.
-From the analysis, point cloud is quite qualified serving as a rough layout guidance in a relatively quick speed, without any further optimization and redundant multi-view images as input.
-The use of point cloud allows our algorithm to be totally training-free and optimization-free, simultaneously being able to produce camera-moving videos quickly with merely one single image or text prompt as input.
+Doubts may arise as to why we do not use a more complex 3D representation which might be more useful.
+Here we take the most prevalent 3D representation–3D Gaussian Splatting  [^6] https://repo-sam.inria.fr/fungraph/3d-gaussian-splatting/, as an example to elaborate on the benefits of using point cloud in three aspects:
+Firstly, concerning the input, point cloud is able to construct the entire scene from one single input image combining techniques of depth estimation and inpainting.
+GS, though also an explicit 3D representation, requires optimization based on images from different views, which means it is neither capable of handling single input image, nor can it leverage 2D inpainting to gradually complete the scene.
+Secondly, from the aspect of time, point cloud can directly lift 2D points onto 3D spaces, while 3DGS demands optimization for each scenario.
+As a training-free approach, our method takes almost no time to generate a video after multi-view images are acquired, but would need more time if 3DGS were applied.
+Lastly, from the task itself, the goal of stage I is not to precisely reconstruct the 3D scene but rather to offer a rough layout guidance.
+In this context, rendered images from point cloud are sufficiently qualified, and no further refinement of the 3D reconstruction is necessary.
+From the analysis, point cloud is quite qualified serving as a rough layout guidance in a relatively quick speed, without any further optimization or redundant multi-view images as input.
+The use of point cloud allows our algorithm to be completely training-free, while still being able to produce camera-moving videos quickly with only one single image or text prompt as input.
 Appendix F Details about Generating 3D Rotation Videos
-3D generation models (Voleti et al., 2024; Melas-Kyriazi et al., 2024; Shi et al., 2023; Chen et al., 2024; Han et al., 2024) are trained on highly-regulated 3D datasets, these datasets are hard to collect, and consist only a limited variety of data types (e.g.
-single static objects with no background).
-As a consequence, 3D generation models exhibit very constrained output distribution and could only produce results in certain styles.
-CamTrol avoid these problems of arduous data collecting, laborious finetuning and output collapse by utilizing the layout prior hidden in video foundation models.
-Not only does it require no training, but this advantage also benefits CamTrol from inheriting most of the prior knowledge inside video foundation models, e.g., the diversity of scenarios, the dynamic of moving objects, temporal consistency, etc.
-Thus CamTrol is able to generate dynamic 3D content, both objects and scenes, in a totally unsupervised and training-free manner, this is what other method cannot achieve yet.
-Compared with regulated datasets, the problem of processing wild pictures in 3D is that some of the parameters is unknown.
+3D generation models (Voleti et al., 2024; Melas-Kyriazi et al., 2024; Shi et al., 2023; Chen et al., 2024; Han et al., 2024) are trained on highly-regulated 3D datasets, these datasets are hard to collect, and include only a limited variety of data types (e.g., single static objects with no background).
+As a consequence, 3D generation models exhibit a very constrained output distribution and can only produce results in certain styles.
+CamTrol avoids these problems of arduous data collection, laborious finetuning, and output collapse by utilizing the layout prior hidden in video foundation models.
+Not only does it require no training, but this advantage also benefits CamTrol from inheriting most of the prior knowledge inside video foundation models, such as the diversity of scenarios, the dynamics of moving objects, temporal consistency, etc.
+Thus, CamTrol is able to generate dynamic 3D content, both objects and scenes, in a totally unsupervised and training-free manner, this is what other methods cannot achieve yet.
+Compared with regulated datasets, the problem of processing wild pictures in 3D is that some of the parameters are unknown.
 In Sec.
-3.1, we’ve mentioned that the camera intrinsic matrix $\mathbf{K}$ and initial extrinsic matrix $\mathbf{P}_{0}$ are set by convention as they’re usually intractable.
-Another crucial parameter concerning 3D video generation is the distance between camera and the content of input image (denote as $f$), note that input image could be synthetic or real.
-Considering most camera rotations are done around the center point, we extract a patch from the very center of the input image, conducting depth estimations on the it and define the distance $f$ as the averaged depth.
-The rotation and transition matrix during camera moving can be formed as :
+3.1, we mentioned that the camera intrinsic matrix $\mathbf{K}$ and the initial extrinsic matrix $\mathbf{P}_{0}$ are set by convention, as they are usually intractable.
+Another crucial parameter concerning 3D video generation is the distance between the camera and the content of input image (denoted as $f$), note that input image could be synthetic or real.
+Considering that most camera rotations are performed around the center point, we extract a patch from the very center of the input image, conduct depth estimations on it, and define the distance $f$ as the average depth.
+The rotation and translation matrices of camera movement can be formed as:
 $\mathcal{R}_{y}$ $=\begin{bmatrix}\cos\theta_{i}&0&-\sin\theta_{i}\\ 0&1&0\\ \sin\theta_{i}&0&\cos\theta_{i}\end{bmatrix},\quad t=\begin{bmatrix}f\sin\theta_{i}\\ 0\\ f-f\cos\theta_{i}\end{bmatrix},$ (6)
 $\text{where}\quad f=\frac{1}{|P|}\sum_{(j,k)\in P}D(x_{0}+j,y_{0}+k).$
-Here $i\in[0,N-1]$ and $\theta_{i}$ refers to rotation angle around $y$ axis at step $i$, $D$ denotes depth estimation of image, $P$ represents the patch around central point $(x_{0},y_{0})$.
-In our experiment, we choose $(j,k)\in[-10,10]$ as the size of patch.
+Here, $i\in[0,N-1]$ and $\theta_{i}$ refer to the rotation angle around the $y$ axis at step $i$, $D$ denotes the depth estimation of the image, and $P$ represents the patch around the central point $(x_{0},y_{0})$.
+In our experiment, we choose $(j,k)\in[-10,10]$ as the size of the patch.
 Appendix G Motion Blur, Problems and Solutions
-Videos produced by CamTrol need to satisfy certain camera movement, and some drastic perspective changes might cause visible trails, recognised as motion blur of objects or scenes.
-This phenomenon will appear to be more indispensable when video base model holds a relatively small generation length (e.g.
-16 frames) as well as the motion scale becomes larger (e.g.
-tilt up for 60 degrees or more).
-To avoid blur issues in controlling video camera motion, we propose several solutions as below:
+Videos produced by CamTrol need to satisfy certain camera movements, and some drastic perspective changes might cause visible trails, recognized as motion blur of objects or scenes.
+This phenomenon will appear to be more indispensable when the video base model has a relatively small generation length (e.g., 16 frames) as well as the motion scale becomes larger (e.g., tilt up by 60 degrees or more).
+To avoid blur issues when controlling video camera motion, we propose several solutions as follows:
   1. 1.
-    According to the analysis above, the blur issue is caused by limited generation frames and large camera movement, thus the most intuitive solution is to either cut down motion scale or utilize a more capable generation model.
-For severe perspective changes, the optimal approach would involve employ video foundation models that support larger generation length (e.g.
-CogVideoX (Yang et al., 2024b) supports generating video with 49 frames).
-This allows model to manage motions of equivalent magnitude while experiencing a smaller moving range between adjacent frames, thereby brings effective alleviation to blur problems.
+    According to the analysis above, the blur issue is caused by limited generation frames and large camera movements, thus the most intuitive solution is to either reduce motion scale or use a more capable generation model.
+For severe perspective changes, the optimal approach would involve employing video foundation models that support larger generation lengths (e.g., CogVideoX (Yang et al., 2024b) supports generating videos with 49 frames).
+This allows the model to manage motions of equivalent magnitude while experiencing a smaller moving range between adjacent frames, thereby bringing effective alleviation to the blur problem.
   2. 2.
-    One can also stack the results of multiple generations to form a complete outcome, i.e., treating the last frame of previous generation as the starting frame for the next, and integrate them as a whole.
-This approach is more suitable when utilizing a image-to-video (I2V) base model.
+    One can also stack the results of multiple generations to form a complete output, i.e., treating the last frame of the previous generation as the starting frame for the next and integrating them as a whole.
+This approach is more suitable when using an image-to-video (I2V) base model.
 Since most open-source video foundation models are text-to-video (T2V), one may consider increasing the step of camera motion inversion $t_{0}$, which guarantees more fidelity towards input images’ content (and motion).
   3. 3.
-    Besides the above two approaches, it is as well a common and convenient choice to apply frame interpolation towards the output.
-Lots of off-the-shelf frame interpolation models are open-sourced and could be found on github.
+    Besides the above two approaches, applying frame interpolation to the output is also a common and convenient choice.
+Many off-the-shelf frame interpolation models are open-source and can be found on GitHub.
   4. 4.
-    Lastly, if video length is not being able to alter, it may be necessary to increase the fps of generated videos for better visualization.
-Although larger fps leads to shorter video duration, it simultaneously makes the visual persistence brought by motions less pronounced, which reduce blur visually.
-In our experiments, we take raw output in all settings.
+    Lastly, if the video length cannot be altered, it may be necessary to increase the fps of the generated videos for better visual quality.
+Although a larger frame rate leads to a shorter video duration, it simultaneously makes the visual persistence brought by motions less pronounced, which reduces blur visually.
+In our experiments, we take the raw output in all settings.